### v1

# FCDM: Sparse-view Sinogram Inpainting with Frequency Domain Convolution Enhanced Diffusion Models

Abstract
Reducing the radiation dose in computed tomography (CT) is crucial, but it often results in sparse-view CT, where the number of available projections is significantly reduced. This reduction in projection data makes it challenging to accurately reconstruct high-quality CT images. In this condition, a sinogram, which is a collection of these projections, becomes incomplete. Sinogram inpainting then becomes essential because it enables accurate image reconstruction with limited projections. Existing models performing well on conventional RGB images for inpainting mostly fail in the case of sinograms. Further, these models usually do not make full use of unique properties, e.g., frequency features and absorption characteristics in the sinogram, and cannot handle large-area masks and complex real-world projections well.
To address these limitations, we propose a novel model called the Frequency Convolution Diffusion Model (FCDM). It employs frequency domain convolutions to extract frequency information from various angles and capture the intricate relationships between these angles, which is essential for high-quality CT reconstruction. We also design a specific loss function based on the unique properties of a sinogram to maintain the consistency in physical properties, which allows the model to learn more effectively even in larger mask areas. We compare FCDM using both simulations and real data with nine inpainting models examples, among which two are designed for sinogram and seven for RGB. The results indicate that our model significantly improves the quality of the inpainted sinograms in terms of both visually and quantitatively, with an SSIM of more than 0.95 and PSNR of more than 30, achieving up to a 33% improvement in SSIM and a 29% improvement in PSNR compared to the baseline.
Introduction
Computed tomography (CT) is a widely used 3D imaging technique for investigating features and phenomena in a broad range of samples and specimens. Advanced CT experiments, such as high-resolution synchrotron-based X-ray CT, can provide spatial resolutions as fine as 20 nm(De Andrade et al. 2016), offering opportunities to observe extremely small features and phenomena in various scientific domains. These include the detection and understanding of defects during battery cycling for the development of durable batteries (Zhao et al. 2024; Müller et al. 2018), mapping the human connectome to study neurodegenerative diseases such as Alzheimer’s and Parkinson’s (Koroshetz et al. 2018; Dyer et al. 2017; Hidayetoğlu et al. 2020), and the development of advanced materials and manufacturing processes (Satapathy et al. 2024). However, the high radiation doses associated with X-ray CT scans can degrade the morphology of features within the imaged samples and pose potential risks for scientific interpretations, making dose reduction a paramount concern (Brenner and Hall 2007).
Figure 1: Schematic of a Synchrotron-Based CT Experiment: The process begins with a sample placed on a rotary stage exposed to X-rays as it rotates, resulting in attenuated photon projections recorded by a detector. These projections are then organized into a 3D dataset, sliced into sinograms for iterative tomographic reconstruction.
As shwon in 1, during a synchrotron-based CT experiment, a sample is placed on a rotary stage and exposed to X-rays as the stage rotates. The transmitted photons lose flux as they propagate through the sample, and these attenuated photons are then recorded by a detector, resulting in X-ray projections. The acquired X-ray projections are organized into a 3D dataset and sliced row-by-row, producing sinograms. Each sinogram is then used to reconstruct the corresponding 3D image slice of the sample, typically using a computationally demanding algorithm (Chen et al. 2019; Hidayetoğlu et al. 2019; Bicer et al. 2017). Low-dose sparse-view CT is the most common approach to lower sample radiation exposure and data acquisition time. However, it typically leads to quality degradation due to the limited number of projections and incomplete sinogram data, which can hinder the reconstructed image quality and constrain the target spatiotemporal resolutions required by the scientific goal.
In recent years, many inpainting models developed for RGB image (Liu et al. 2024; Zhang et al. 2023; Ko and Kim 2023) have shown success in filling missing data or pixels of RGB images, but their application to sinograms, the raw projection data in CT imaging, is not efficient. Sinograms possess unique structural and statistical properties that are distinct from RGB images, requiring specialized approaches for effective inpainting. RGB inpainting models are primarily designed for images where light reflects off objects, forming distinct, non-overlapping features. In contrast, X-ray images, such as sinograms, involve overlapping features due to the penetrating nature of X-rays, making them fundamentally different from RGB images.
Some recent models designed for sinogram inpainting (Yao et al. 2024; Valat, Farrahi, and Blumensath 2023a, b; Xie et al. 2022) treat sinograms merely as regular images without leveraging their unique characteristics. These models often struggle to handle larger mask ratios and also tend to falter when applied to more complex real-world data. Other works (Li et al. 2019a; Wagner et al. 2023; Li et al. 2019c) attempt to incorporate certain sinogram-specific features by modifying the loss function terms, such as a term that checks the periodicity in the inpainted sinogram, which refers to the repeating patterns that occur due to the circular nature of CT scans. However, this characteristic is not always present, particularly in cases of incomplete scans or non-uniform sampling.
More recently, Diffusion Models (Sohl-Dickstein et al. 2015; Rombach et al. 2022) have shown promising results in inpainting RGB images. However, their performance in sinogram inpainting is less satisfactory due to their failure to leverage the inherent (physical) properties of sinograms due to, e.g., data acquisition patterns.
Within this context, this work introduces a novel diffusion-based inpainting model (called FCDM) designed specifically for sinograms. To capture the physical properties inherent in sinograms, such as material absorption rates and density variations, while also simplifying complex convolution operations into element-wise multiplications. FCDM leverages the frequency domain information of sinograms through frequency domain convolutions. This enables the extraction of the frequency domain features from various projection angles and also the relationships between various angles. Additionally, to ensure the physical accuracy and preserve the frequency characteristics of the sinogram, we propose a custom loss function that aligns with the specific properties of sinograms, enhancing the model’s inpainting quality. Specifically, this work consists of three main contributions as follows:
  • A novel approach is specifically designed for sinogram data by applying frequency domain convolutions. This allows the model to effectively capture and correlate frequency features from different projection angles and their interrelationships.
  • A new loss function is developed to align with the unique statistical properties of sinograms. This loss function is closely integrated with the frequency domain convolution, ensuring that the frequency characteristics extracted during convolution are accurately reflected and preserved, further enhancing the inpainting quality.
  • An extensive evaluation is conducted on both a real-world dataset (De Carlo et al. 2018), and two common synthetic and simulated datasets. Evaluation results demonstrate that our model (FCDM) for the first time achieves high accuracy for the complex real-world dataset. The subsequent ablation study proves the efficacy of our new designs, frequency domain convolutions, and the new loss function.
More specifically, our experimental results demonstrate that FCDM significantly improves the quality of sinogram inpainting, outperforming existing methods in both visual assessments and quantitative metrics, with more than 0.95 in the SSIM (Wang et al. 2004) metric and more than 30 in the PSRN metric, achieving an improvement up to 33% in SSIM and 29% in PSNR. The model can handle larger missing areas (or a limited number of projections) and complex real-world sinograms compared to the existing solutions, showcasing its robustness and adaptability.
Related Work and Our Opportunities
Several traditional mathematical methods have been employed for sinogram inpainting. Interpolation methods, such as linear interpolation (Herman and Meyer 1993) and spline interpolation (Unser 1999), estimate the missing sinogram data based on the known data. These methods are computationally efficient and easy to implement but often produce suboptimal results when dealing with large missing areas. Total Variation (TV) Minimization (Chambolle and Lions 1997) which is originally developed for image denoising and reconstruction has also been applied to sinogram inpainting. It works by promoting sparsity in the gradient domain, making it effective for inpainting sinograms with small missing regions, but struggle with larger gaps. Although, these techniques enhance the image quality, they ignore the physical properties of the data acquisition.
More recently, deep learning-based methods have shown great promise in the domain of RGB image inpainting. Convolutional Neural Networks (CNNs) (Xie, Xu, and Chen 2012), U-Net (Ronneberger, Fischer, and Brox 2015), and Generative Adversarial Networks (GANs) (Yu et al. 2018) have been applied to enhance image inpainting, providing significant improvements in inpainting quality. These methods have been particularly effective for RGB images, leveraging spatial and contextual information to fill in missing regions. In addition to these methods, Diffusion Models (Sohl-Dickstein et al. 2015; Rombach et al. 2022) have emerged as a powerful tool for image inpainting and in particular, have shown excellent performance in the inpainting of RGB images. These models leverage a diffusion process to iteratively refine image features, achieving high-quality inpainting results. However, when applied to sinograms, these models fall short because they are designed with the assumptions and properties of RGB data, neglecting the unique characteristics of sinogram data.
Several deep learning-based models have been specifically proposed for sinogram inpainting. (Yao et al. 2024; Zhao et al. 2018) utilize U-Net to directly interpolate missing sinogram data, while (Valat, Farrahi, and Blumensath 2023a; Li et al. 2019b; Valat, Farrahi, and Blumensath 2023b; Xie et al. 2022) employ GANs to generate realistic sinogram patches. These models typically treat the sinograms as normal RGB images and apply deep learning techniques without significant modifications to accommodate the specific characteristics of sinogram data. This lack of customization results in suboptimal performance, especially when dealing with large-area masks. Other deep learning-based models apply unique characteristics of sinograms to tailor the model to fit in sinogram inpainting. (Li et al. 2019a, c) incorporate periodicity into the loss function to leverage the cyclical nature of certain sinograms. However, this approach is not applicable in cases of incomplete scans or non-uniform sampling. (Wagner et al. 2023; Li et al. 2019b) compare the reconstructed images from the inpainted and ground truth sinograms in their loss calculations, while significantly increasing computational overhead, leading to longer training times.
Frequency domain convolutions such as Fast Fourier Convolution (FFC) (Chi, Jiang, and Mu 2020) can efficiently capture global features and extract features in the frequency domain. Traditional FFC methods can extract global features utilizing a small amount of computation resources compared to traditional convolutions, which can improve the inpainting quality and performance without too much computational overhead. However, FFC is still designed for RGB image inpainting and does not adequately address the unique frequency characteristics of sinograms or their complex inter-angle relationships.
FCDM Design and Analysis
(a) Overview of FCDM: In stage 1, the Encoder-Decoder network is trained to generate a latent representation capturing both spatial and frequency features. In stage 2, with fixed parameters from stage 1, random masking information is embedded using the K, Q, V formulation.
(b) Overview of Freq Conv: After obtaining the latent representation from the Encoder, both standard convolution and frequency domain convolution are applied to the latent features. The results from these two operations are then concatenated. Finally, a subsequent standard convolution is performed to restore the original number of channels.
Figure 2: Overview of FCDM
FCDM is a two-stage process model that first employs an Encoder-Decoder network to capture both spatial and frequency domain features to better represent the full image by a latent, followed by a Diffusion process that takes mask information and then does inpainting work. Figure 2 shows a brief overview of it. In the first stage, the Encoder encodes the input sinogram into a latent representation, which is then enhanced using frequency domain convolutions. This enriched latent representation is subsequently decoded to recovef the sinogram. Once the training of this stage is finished, the Encoder-Decoder network’s parameters are fixed and does not affect the training process of the second stage. The second stage involves training a Diffusion Model on the latent space, where various random masks are applied to simulate real-world conditions with missing or corrupted data. This allows the model to effectively handle complex sinogram inpainting scenarios.
Frequency Domain Convolution
FCDM begins with the input sinogram being encoded into a latent representation, which is then enhanced through frequency domain operations, and finally decoded back into the reconstructed sinogram.
The Encoder $\mathcal{E}(\cdot)$ transforms the input sinogram $s\in{\mathbb{R}}^{C\times H\times W}$ into a latent representation $h\in{\mathbb{R}}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$:
$h^{C^{\prime}\times H^{\prime}\times W^{\prime}}=\mathcal{E}(s^{C\times H\times W}),$ (1)
where $C$, $H$, and $W$ represent the number of channels, height, and width, respectively. Once the latent representation $h\in{\mathbb{R}}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$ is obtained, we perform a Real Fourier Transform (RFFT) (Cooley and Tukey 1965) along the width dimension on the latent representation to convert it into the frequency domain. The output $\hat{h}_{w}$ is a complex-valued tensor:
$\hat{h}_{w}=\mathcal{F}_{w}(h)=\Re(\hat{h}_{w})+j\Im(\hat{h}_{w}).$ (2)
Next, we reshape the complex tensor $\hat{h}_{w}$ into a real-valued tensor by concatenating the real and imaginary components along the channel dimension:
${\hat{h}_{w}}^{C^{\prime}\times H^{\prime}\times W^{\prime}}\longrightarrow{\tilde{h}_{w}}^{2C^{\prime}\times H^{\prime}\times\frac{W^{\prime}}{2}}.$ (3)
This operation doubles the number of channels, allowing the model to perform convolution on both components.
In the spatial domain, convolution operations are computationally expensive due to the need to sum over sliding windows of input data. However, by transforming the convolution operation into the frequency domain, we can exploit the Convolution Theorem, which states that convolution in the spatial domain corresponds to Hadamard multiplication in the frequency domain:
$\hat{y}_{w}=\mathcal{G}(\mathcal{K}_{w}\cdot{\tilde{h}_{w}})$ (4)
$\mathcal{G}$ is a non-linear activation function, $\cdot$ denotes Hadamard multiplication, and $\mathcal{K}_{w}$ is the Fourier-transformed convolution kernel. This Hadamard multiplication is computationally more efficient than traditional spatial domain convolution because it avoids the need for sliding windows and reduces the operation to simple multiplications.
After convolution, the result is reshaped back into a complex-valued tensor by reconstituting the real and imaginary parts:
$\hat{y}_{w}=\tilde{y}_{w,\Re}+j\tilde{y}_{w,\Im}$ (5)
The complex tensor $\hat{y}_{w}$ is then transformed back into the spatial domain using the Inverse Real Fourier Transform (Cooley and Tukey 1965):
$y_{w}=\mathcal{F}_{w}^{-1}(\hat{y}_{w})$ (6)
This yields the spatial domain representation $y_{w}$, and similarly, a process is applied along the height dimension and gets $y_{h}$.
The outputs from the frequency domain convolutions along the width and height dimensions are combined:
$y={\alpha}_{w}y_{w}+{\alpha}_{h}\beta y_{h}$ (7)
The contribution from $y_{w}$ emphasizes the frequency characteristics of the sinogram across different angles, capturing essential features within each projection. Conversely, $y_{h}$ focuses on the relationships between different angles, encapsulating the global structure and dependencies across the entire sinogram. The coefficients ${\alpha}_{w}$ and ${\alpha}_{h}$ are selected to balance these two components, ensuring that the model captures both localized frequency information and the overarching global context.
Once the enhanced latent representation $y$ is obtained, it is passed through the Decoder $\mathcal{D}(\cdot)$:
$\overline{s}^{C\times H\times W}=\mathcal{D}(h^{C^{\prime}\times H^{\prime}\times W^{\prime}})$ (8)
This decoding process translates the frequency-enhanced latent features back into the spatial domain, yielding the final inpainted sinogram $\overline{s}\in{\mathbb{R}}^{C\times H\times W}$.
The Encoder and Decoder remain fixed during the second stage of training, which not only reduces the computational burden but also facilitates easy adaptation for different downstream tasks, such as full CT reconstruction. This design choice enhances the model’s versatility and efficiency in diverse applications.
Comparison Related to Convolution Locations
There are two strategies for applying frequency domain convolutions in sinogram inpainting: i) performing convolution after downsampling the original sinogram and ii) performing frequency domain convolution within the latent space, post-encoding. The second strategy processes the full-resolution input sinogram through the Encoder, which allows for capturing and preserving a richer and more comprehensive set of features since Encoder captures global features better than downsampling. It enables the model to better capture detailed features from each angle as well as the intricate relationships between various angles and results in better overall inpainting quality compared to the first strategy. More details of ablation evaluation results are shown in Ablation Studies. Analysis of complexity of the two strategies is in the supplementary materials.
Tailored Loss Functions
In designing the loss function for our model, we integrate various established techniques to optimize both perceptual quality and structural accuracy. Pixel loss measures pixel-wise differences to ensure accurate reconstruction. Perceptual loss (Zhang et al. 2018) focuses on maintaining perceptual similarity by comparing high-level features. Adversarial loss (Goodfellow et al. 2014; Lim and Ye 2017) uses a discriminator to encourage the generation of realistic images. Codebook loss (Van Den Oord, Vinyals et al. 2017) prevents collapse in the VQ-VAE’s latent codebook by maintaining diversity. Adaptive GAN loss (Karras, Laine, and Aila 2019) adjusts the adversarial loss weight dynamically, stabilizing the training process.
While these loss functions are highly effective for RGB image inpainting, their direct application to sinogram data is less successful. RGB images are spatially structured and benefit from perceptual and pixel-level losses that align well with human visual perception. In contrast, sinograms represent projection data, which encodes information across different angles and requires consideration of the underlying physical properties (Kak and Slaney 2001).
Total Projection Consistency Loss
In the context of CT, one essential physical property is the relationship between the sum of projection data across all angles and the total absorption of the object being imaged. The frequency domain convolution extracts frequency features that encapsulate the physical characteristics of the object, which are inherently linked to its absorption behavior. We first tailor the loss function to enforce this consistency with the total absorption characteristic inherent in CT imaging. By incorporating this term, we ensure that the model not only captures the detailed frequency features but also maintains accurate physical consistency in the inpainted sinogram and further reconstructed images.
The sum of the projection data across all angles is ideally equals the total absorption of the object, i.e.:
[…]
where $P_{\theta}(s)$ represents the projection data at angle $\theta$ and position $s$ on the detector, and $\mu(x,y)$ is the linear attenuation coefficient at the point $(x,y)$ in the object.
To enforce this consistency within the loss function, we introduce the total projection consistency loss $\mathcal{L}_{\text{absorp\_sum}}$:
[…]
This loss term ensures that the reconstructed image maintains the physical principle where the total detected radiation equals the total absorption of the object.
The second component of $\mathcal{L}_{\text{absorp\_sum}}$ involves calculating the total absorption of the object using the inverse Radon transform (iRadon transform) (Radon 1917),
$\mu(x,y)=\mathcal{R}^{-1}\{P_{\theta}(s)\}.$ (11)
Frequency Domain Consistency Loss
The frequency domain loss term directly evaluates the accuracy of the frequency information extracted by the frequency domain convolutional layers. By minimizing the difference between the frequency representations of the inpainted and original sinograms, this loss ensures that the frequency domain convolution effectively captures and preserves the essential frequency characteristics.
To calculate the difference in the frequency domain, we first need to transform the sinograms from the spatial domain to the frequency domain. This transformation is accomplished using the 2D Fourier Transform (Cooley and Tukey 1965). For the inpainted sinogram $\hat{\mu}(u,v)$ and the ground truth $\mu(u,v)$, their frequency domain representations $\hat{F}(u,v)$ and $F(u,v)$ are given by:
$\overline{F}(u,v)=\mathcal{F}\{\overline{\mu}(x,y)\}.$ (12)
$F(u,v)=\mathcal{F}\{{\mu}(x,y)\}$ (13)
To quantify this difference and incorporate it into the loss function, we calculate the sum of L2 norm in the frequency domain:
$\mathcal{L}_{\text{freq}}=\sum_{u}\sum_{v}\left\{\overline{F}(u,v)-F(u,v)\right\}^{2}.$ (14)
The final loss function is a weighted combination of several terms, each carefully designed to enforce different aspects of the inpainting process. By integrating multiple loss components, the loss can better leverage the unique characteristics of sinograms, which makes the loss more fit for sinogram inpainting and effectively handles the complex nature of sinogram data.
Mask Generation and Training
In the second stage of our method, we introduce a masking strategy within the Diffusion process to enhance the model’s robustness and ability to handle complex inpainting conditions. Specifically, we automatically generate masks of varying sizes and locations, randomly masking different angles of the sinogram, which simulates the real-world scenario where certain parts of the sinogram data may be missing or corrupted.
To effectively integrate these masks into the Diffusion process, we employ a Transformer-based (Vaswani et al. 2017) mechanism, where the mask information is embedded using the Key (K), Query (Q), and Value (V) formulation. By leveraging the attention mechanism, the model can focus on relevant features and better infer missing information from the available data.
Evaluation
This section evaluates FCDM’s performance on three datasets: Real-world (De Carlo et al. 2018), Shapes generated using (Van der Walt et al. 2014), and Shepp2d generated using (Gürsoy et al. 2014), and compares it with nine baselines, in which UsiNet (Yao et al. 2024) and SionTx (Liu, Kettimuthu, and Foster 2022) are designed for sinogram inpainting, and StrDiffusion (Liu et al. 2024), CoPaint (Zhang et al. 2023), CMT (Ko and Kim 2023), hourglass (Deng et al. 2022), RePaint (Lugmayr et al. 2022), MISF (Li et al. 2022), and LaMa (Suvorov et al. 2022) are for more general image inpainting. This evaluation has three main objectives: 1) demonstrating FCDM outperforms all baselines, specifically achieving high accuracy on the complex real-world dataset for the first time (Sec. Accuracy Comparisons with Baselines); 2) studying the underlying reasons for performance gains with an ablation study, proving the efficacy of our two key designs (Sec. Ablation Studies); and 3) studying FCDM’s performance with a key factor in practice, mask ratio, showing its robustness (Sec. Mask Ratio Analysis).
Experimental Setup
(a) Mask
(b) Ground Truth
(c) Ours
(d) UsiNet
(e) SinoTx
(f) StrDiffusion
(g) Copaint
Figure 3: Visual Results: Each line shows an visual result of the Real-world dataset. Details of visual results for Shapes and Shepp2d datasets are in supplementary materials.
Implementation Details
We train our model using the Polaris supercomputer at Argonne Leadership Computing Facility (ALCF) resources at Argonne National Laboratory (ANL) [^1] Polaris Cluster at ALCF: https://www.alcf.anl.gov/polaris. Polaris supercomputer consists of 560 compute nodes, each of which is equipped with 4 NVIDIA A100 GPUs (connected via NVLink). We use PyTorch version 2.1.0 and CUDA version 12.2. The dimensions of the input images are 512x512. For frequency domain convolutions, the weights for the $H$ and $W$ dimensions are adjusted to 0.55 and 0.45, respectively. The loss function is a weighted combination of several components with the following coefficient: pixel loss $\mathcal{L}_{pixel}=1.0$, perceptual loss $\mathcal{L}_{perceptual}=0.5$, adversarial loss $\mathcal{L}_{adv}=0.05$, codebook loss $\mathcal{L}_{cb}=0.5$, absorption consistency loss $\mathcal{L}_{absorp\_sum}=0.1$, and frequency domain loss $\mathcal{L}_{freq}=0.1$. During training, the mask ratio is randomly chosen within a range from $0.1$ to $0.9$ with the corresponding proportion of angles being randomly masked. The learning rate is adjusted based on the formula: $lr={lr}_{base}\times batch\,size\times GPU\,num\times grad\,accum\,steps$.
Dataset
Three datasets are used in our experiments, each containing 100k samples with a resolution of $512\times 512$:
Real-world dataset (De Carlo et al. 2018) consists of real sinogram images obtained from various materials and objects, derived from actual synchrotron radiation CT scans, including Advanced Photon Source (APS [^2] Advanced Photon Source: https://www.aps.anl.gov) at ANL. This dataset, provided by the x-ray tomography data bank, includes a range of noise levels and artifacts typical in real-world imaging, making it significantly more complex than the simulated and synthetic datasets. This real-world dataset is particularly challenging due to the inclusion of dynamic (Mohan et al. 2015) and in situ (Pelt and Batenburg 2013) systems, which are representative of materials imaged at APS and other synchrotron facilities. This complexity provides a rigorous test of the model’s robustness and its ability to handle practical challenges in sinogram inpainting. Shape dataset generated using the scikit-image library (Van der Walt et al. 2014) contains a variety of simulated geometric shapes, including circles, rectangles, and triangles. This dataset introduces a range of geometric complexities, allowing us to assess the model’s ability to reconstruct shapes with sharp edges and varying curvature. Shepp2d dataset generated using the Tomoly library (Gürsoy et al. 2014) comprises simulated shepp-logan phantoms, which offer a controlled environment for evaluating the accuracy and stability of reconstruction algorithms. These phantoms are less complex, making them ideal for testing basic inpainting capabilities.
Metrics
This evaluation leverages two metrics, Structural Similarity Index Measure (SSIM) (Wang et al. 2004) and Peak Signal-to-Noise Ratio (PSNR), to quantitatively assess the quality of the inpainted sinograms. SSIM assesses structural similarity by considering changes in structure, luminance, and contrast, making it effective for evaluating how well the inpainted sinogram preserves the structural integrity necessary for sinogram inpainting. PSNR measures the ratio between the maximum possible signal power and the noise affecting image fidelity, providing a quantitative assessment of how closely the inpainted sinogram approximates the ground truth, which is crucial for minimizing artifacts.
Accuracy Comparisons with Baselines
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.961 0.965 0.954 32.3 31.4 31.4
UsiNet 0.460 0.482 0.494 12.8 14.5 14.9
SionTx 0.572 0.597 0.643 20.4 21.7 21.9
StrDiffusion 0.462 0.508 0.507 13.7 15.7 16.2
CoPaint 0.725 0.738 0.761 23.0 23.2 23.9
CMT 0.436 0.440 0.453 16.1 16.2 16.1
hourglass 0.592 0.601 0.628 21.3 21.7 22.7
RePaint 0.720 0.754 0.766 24.6 25.2 26.3
MISF 0.706 0.717 0.723 25.1 25.2 26.4
LaMa 0.706 0.716 0.715 22.1 22.9 24.6
Table 1: Quantitative Accuracy Results: FCDMis compared with all 9 baselines on the Real-world dataset. The comparison results of Shape and Shepp2d datasets are in the supplementary materials.
This section evaluates FCDM’s performance by comparing it against all nine baselines on three datasets. For each comparison, we focus on evaluating the SSIM and PSNR metrics specifically within the masked regions, rather than across the entire image. This approach allows us to precisely measure the model’s ability to reconstruct the most challenging portions of the sinogram, where data is missing or corrupted.
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.961 0.965 0.954 32.3 31.4 31.4
w/o Freq Conv 0.789 0.816 0.813 26.7 27.0 27.3
w/o $H$ Conv 0.846 0.847 0.846 26.1 26.5 27.5
w/o $W$ Conv 0.880 0.884 0.891 27.1 27.9 28.2
Table 2: Ablation Results Related to Freq Conv: The experimental results show that the model’s performance is greatly impacted by the frequency domain convolution. Furthermore, there is a marginally more noticeable drop in performance when the $H$ dimension convolution is absent than when the $W$ dimension convolution is absent. $W$ dimension convolution retrieves important frequency information inside each individual angle, whereas the $H$ dimension frequency convolution captures the inter-angle correlations.
Table 1 summarizes the performance of our model and the baseline models on Real-world dataset [^3] Due to space limitation, please find our results on Shape and Shepp2d datasets in our supplementary materials.. In addition to the quantitative analysis, we also provide visual comparisons between the outputs of our model and the baseline models. Figure 3 illustrates the inpainted sinograms when the mask ratio is equal to $0.8$, focusing on showing the visual quality and structural fidelity of the inpainted sinograms. Our experimental results show that our model significantly outperforms all baselines on the more complex Real-world dataset. The simplicity of the Shapes and Shepp2d datasets allows previous baselines to achieve good results, making our model’s improvements relatively modest (but still obvious) in these cases. Real-world dataset presents greater challenges, and our model excels by effectively leveraging the unique properties of sinogram, leading to a clear and substantial advantage in inpainting quality over the other methods.
Ablation Studies
Our ablation studies are divided into two key sections, each designed to validate the contributions of a component of our model, frequency domain convolution and tailored loss functions. More specifically, to assess the impact of frequency domain convolutions, we conducted experiments where this component was either disabled or modified.
Impact of Frequency Domain Convolutions
First, we compare the performance of the model with and without frequency domain convolutions and without $H$ or $W$ dimension convolutions. Results in Table 2 clearly show that the inclusion of frequency domain convolutions significantly enhances the model’s ability to inpaint sinograms with greater accuracy, particularly in challenging scenarios with significant amount of missing data. More details of visual results are in supplementary materials.
Locations SSIM PSNR
Ds Followed by Freq Conv 0.962 33.9
Latent Freq Conv 0.987 35.9
Table 3: Ablation Results Related to Freq Conv Locations: The experiments are done only on the first stage. The results show that, conducting frequency domain convolutions on the latent after the input has been passed through an Encoder produces better performances.
Additionally, we evaluate the locations of the frequency domain convolutions within the model, comparing their application before and after the Encoder. This comparison focuses on the Encoder-Decoder component. Its results in Table 3 indicate that applying frequency domain convolutions after the Encoder, within the latent space, leads to better inpainting quality. This is likely due to the better abstraction and compression ability of Encoder compared to downsampling, which allows the frequency domain operations to better capture features from various angles.
Contribution of Tailored Loss Functions
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.961 0.965 0.954 32.3 31.4 31.4
w/o $L_{absorp\_sum}$ 0.917 0.930 0.927 29.0 29.4 29.6
w/o $L_{freq}$ 0.923 0.935 0.926 29.6 29.4 29.8
Table 4: Ablation Results Related to Tailored Loss: The results illustrate the model’s performance on the real-world dataset when specific terms are removed from the loss function. It is evident that the model’s performance degrades when either $L_{absorp\_sum}$ or $L_{freq}$ is excluded.
The second part of our ablation study examines the role of the tailored loss functions, specifically the absorption consistency loss and the frequency domain loss. We conduct experiments where each of these losses is individually removed from the training process. Results in Table 4 demonstrate that the tailored losses are essential for achieving high-quality reconstructions. Removing the absorption consistency loss results in a failure to fully adhere to the physical properties of CT imaging. Similarly, the absence of the frequency domain loss results in a degradation in the ability to maintain the correct frequency characteristics, leading to less accurate inpainting.
Mask Ratio Analysis
Figure 4: Relationships between Mask Ratio and Metrics: As mask ratios increase, SSIM and PSNR tend to rise but are not strictly monotonic, exhibiting fluctuations.
During our experiments, we observe an intriguing trend: as the mask ratio increases, the SSIM and PSNR metrics in the masked regions of the inpainted sinograms also tend to improve. Figure 4 shows the plot of this trend of the real-world dataset. We believe this phenomenon can be attributed to the model’s ability to better capture and learn global features when faced with larger masked areas. With a higher mask ratio, the model is less likely to overfit to local features.
Conclusion and Future Work
In this work, we introduced a novel model FCDM for sinogram inpainting that applies frequency domain convolutions and tailored loss functions. We demonstrate that our model excels on real-world data and performs competitively on synthetic data through experiments. Ablation studies further validate the importance of each component, including frequency domain processing and specialized loss functions. This work lays the groundwork for extending our approach to complete CT reconstruction, where the model could be further developed to take sinograms as input and directly produce fully reconstructed CT images.
Appendix A Complexity Analysis of Different Convolution Locations
Downsampling Followed by Frequency Domain Convolution
Downsampling reduces the spatial dimensions of the input image from $H\times W$ to $\frac{H}{r}\times\frac{W}{r}$, where $r$ is the downsampling factor:
$\mathcal{O}_{down}=\mathcal{O}\left(\frac{2CWH}{r^{2}}\right).$ (15)
After downsampling, half of the channels undergoes standard convolution, with complexity:
$\mathcal{O}_{s}=\mathcal{O}\left(C^{2}k_{s}^{2}\frac{WH}{r^{2}}\right),$ (16)
where $k_{s}$ represents the size of the convolution kernel.
For the other channels, a Fourier Transform is applied, followed by a frequency domain convolution, which is equal to Hadamard multiplication in the spatial domain, and then an inverse Fourier Transform:
$\begin{split}\mathcal{O}_{f}&=\mathcal{O}\left(C\frac{WH}{r^{2}}\log{\frac{WH}{r^{2}}}\right)+\mathcal{O}\left(C\frac{WH}{r^{2}}\right).\\ \end{split}$ (17)
Then a convolution is performed to restore the channel count:
$\mathcal{O}_{r}=\mathcal{O}\left(2C^{2}k_{r}^{2}\frac{WH}{r^{2}}\right)$ (18)
Finally, the result after the restore convolution is passed through an Encoder $\mathcal{E}(\cdot)$. The complexity of it depends on the number of layers $L$, the number of channels $C_{i}$ in each layer, and the size of the convolutional kernels $k_{i}$:
$\mathcal{O}_{e}=\mathcal{O}\left(\sum_{i=1}^{L}C_{i-1}C_{i}k_{i}^{2}\frac{H_{i}W_{i}}{r^{2}}\right),$ (19)
where $H_{i}$ and $W_{i}$ are the dimensions of the feature map at that layer.
Latent Space Frequency Domain Convolution
The input image is first passed into an Encoder $\mathcal{E}(\cdot)$, which transforms it into a latent representation $h$ with dimensions $C_{L}\times H_{L}\times W_{L}$:
$\mathcal{O}_{e}=\mathcal{O}\left(\sum_{i=1}^{L}C_{i-1}C_{i}k_{i}^{2}H_{i}W_{i}\right).$ (20)
After encoding, standard convolution is applied to the latent:
$\mathcal{O}_{s}=\mathcal{O}(C_{L}^{2}k_{s}^{2}H_{L}W_{L}).$ (21)
Frequency domain convolution followed by an inverse Fourier Transform are also applied:
$\begin{split}\mathcal{O}_{f}&=\mathcal{O}(C_{L}H_{L}W_{L}\log{H_{L}W_{L}})+\mathcal{O}(C_{L}H_{L}W_{L}).\\ \end{split}$ (22)
Finally, a convolution to restore the channel count is performed:
$\mathcal{O}_{r}=\mathcal{O}(2C_{L}Ck_{r}^{2}H_{L}W_{L})$ (23)
Appendix B Supplementary Evaluation Results
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.965 0.977 0.977 36.2 36.3 36.6
UsiNet 0.509 0.549 0.546 14.1 13.7 15.3
SionTx 0.746 0.743 0.796 26.5 26.9 27.1
StrDiffusion 0.524 0.552 0.600 14.3 15.1 15.9
CoPaint 0.903 0.929 0.908 30.2 31.0 33.0
CMT 0.454 0.469 0.477 19.9 20.2 21.0
hourglass 0.807 0.813 0.823 27.1 27.5 28.2
RePaint 0.915 0.918 0.926 33.1 33.6 33.69
MISF 0.716 0.739 0.753 26.4 27.6 27.8
LaMa 0.747 0.762 0.779 26.1 27.8 28.6
(a) Geometric Shapes Dataset
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.966 0.978 0.977 37.7 37.2 37.8
UsiNet 0.494 0.499 0.492 13.1 15.0 15.6
SionTx 0.967 0.968 0.962 37.2 37.5 36.3
StrDiffusion 0.516 0.554 0.562 15.5 16.3 16.7
CoPaint 0.932 0.947 0.946 32.0 31.8 32.7
CMT 0.504 0.507 0.518 20.8 22.6 22.7
hourglass 0.826 0.843 0.851 27.5 27.7 28.1
RePaint 0.912 0.931 0.934 33.2 34.3 35.2
MISF 0.731 0.749 0.743 25.8 26.5 26.9
LaMa 0.698 0.708 0.720 25.0 25.8 26.6
(b) Simulated Shepp-logan Phantom Dataset
Table 5: Results of Shape and Shepp2d Datasets: The simplicity of these two datasets allows previous baselines to achieve good results but our model still has comparable performances.
(a) Mask
(b) Ground Truth
(c) UsiNet
(d) SinoTx
(e) StrDiffusion
(f) CoPaint
(g) Ours
Figure 5: Visual Results of Shape and Shepp2d Datasets: Line 1 is a visual result of Shape dataset and line 2 is a visual result of Shepp2d datset.
Ours
W/o Freq Conv
Figure 6: Visual Ablation Results: Results without frequency domain convolution are blurred.
Table 5 and Figure 5 shows the results of Shape and Shepp2d datasets with other baselines. Our model’s improvements are relatively modest but still comparable to existing baselines. Figure 6 shows the visual results of ablation studies of Real-world dataset, which better proves the impact of frequency domain convolutions.