### v2

# FCDM: Sparse-view Sinogram Inpainting with Frequency Domain Convolution Enhanced Diffusion Models

Abstract
Computed tomography (CT) is an imaging technique that uses X-ray projections from multiple rotation angles to create detailed cross-sectional images, widely used in industrial inspection and medical diagnostics. Reducing the projection data in CT scans is often necessary to decrease radiation exposure, scanning time, and computational costs. However, this reduction makes accurate image reconstruction challenging due to the incomplete sinogram. Existing RGB inpainting models struggle with severe feature overlap, while current sinogram-specific models fail to employ efficient feature extraction methods that account for the physical principles underlying the sinogram generation process. To tackle these challenges, we introduce the Frequency Convolution Diffusion Model (FCDM), a novel diffusion-based inpainting framework tailored for sinogram data. FCDM leverages frequency-domain convolutions to capture global and fine-grained structural features, effectively disentangling overlapping components across projection angles. Additionally, we propose a custom loss function that incorporates unique sinogram properties of total absorption consistency and frequency-domain consistency. Extensive experiments on synthetic and real-world datasets demonstrate that FCDM significantly outperforms existing methods, achieving SSIM over 0.95 and PSNR above 30 dB, with improvements of up to 33% in SSIM and 29% in PSNR compared to baselines.
1 Introduction
X-ray computed tomography (XCT) is a fundamental 3D imaging technique used at synchrotron radiation facilities for investigating features and phenomena in a broad range of samples and specimens. For instance, the nano-CT beamlines at the Advanced Photon Source (APS) at Argonne National Laboratory (ANL) provide cutting-edge imaging capabilities, achieving sub-20 nm spatial resolution. These experiments enable detailed studies critical to various fields, such as identifying and analyzing defects during battery cycling to support the design of long-lasting batteries [50, 30], mapping the human connectome to investigate neurodegenerative diseases like Alzheimer’s and Parkinson’s [20, 11, 16], and advancing material science and manufacturing techniques [34]. In synchrotron-based CT, a sample is placed on a rotary stage and exposed to collimated X-rays, with attenuated photons captured by a detector as parallel beam projections. These projections are organized into a 3D dataset and sliced row-by-row, producing sinograms, which are then used to reconstruct 3D image slices using computationally intensive algorithms [4, 15, 1].
Figure 1: High-level design of FCDM. More details are in Sec. 3
Although XCT scans provide high precision, the high radiation dose can degrade sample features, and increase radiation exposure, scanning time, and computational costs, making dose reduction essential [2]. Low-dose sparse-view CT is the most common approach to lower sample radiation exposure and data acquisition time. However, it typically leads to quality degradation due to incomplete sinogram data, which can hinder the reconstructed image quality and constrain the target spatiotemporal resolutions required by the scientific goal. Consequently, there is a critical need for effective inpainting techniques for sparse sinogram data and further enhancing the quality of reconstructed images.
In recent years, many inpainting models developed for RGB images [26, 48, 19, 9, 28, 21, 37] have demonstrated strong performance in completing missing data or pixels. However, the distinct characteristics of sinogram and RGB images necessitate approaches tailored to the unique properties of sinogram data. This necessity is driven by three fundamental factors. First, RGB images and synchrotron-based CT sinograms are fundamentally different in their imaging mechanisms and resultant features. RGB images capture surface light reflections with clear spatial boundaries, making feature overlap minimal. In contrast, sinograms are generated from cumulative X-ray projections through entire objects, which leads to significant feature overlap [8]. A single pixel in a sinogram may encode information from multiple internal structures (e.g., edges of dense materials, voids due to internal defects, and variations in layered material densities). The inpainting errors in sinograms can propagate through the CT reconstruction process, leading to significant artifacts in reconstructions. Second, traditional CNN-based inpainting models are optimized for natural images with sharp edges and fine textures. In sinograms, low-frequency components represent the overall density distribution, while mid- and high-frequency components capture intricate patterns, such as periodic variations or internal defect signatures. Spatial convolutions in CNNs struggle to disentangle these mixed features, as they are not designed to separate information based on frequency. Third, existing sinogram inpainting methods typically lack specialized feature extraction strategies for the cumulative nature of sinograms. Traditional interpolation and regularization techniques [14, 38, 3] often fail to handle large missing regions and overlook essential physical properties. Although recent machine learning approaches [47, 39, 40, 46] have been applied to sinogram inpainting, they generally treat sinograms as standard RGB images, disregarding their unique physical and structural properties. Other works [22, 44, 24, 36] attempt to integrate sinogram-specific characteristics into machine learning models, yet they do not utilize efficient feature extraction methods that account for the physical principles underlying the sinogram generation process, resulting in suboptimal performance.
Given this context, we introduce a novel frequency-domain diffusion-based inpainting model, termed FCDM, designed specifically for XCT sinograms. Our approach utilizes frequency transforms to separate overlapping components, enabling precise extraction of structural information essential for accurate sinogram inpainting. Additionally, in iteratively refining complex data distributions, FCDM leverages diffusion models’ strengths [35, 33]. Unlike conventional models that overlook sinograms’ physical properties, and existing sinogram-specific machine learning models lacking targeted feature extraction, FCDM employs frequency-domain convolutions to extract distinct object features across multiple projection angles. Low frequencies capture global structures, while high frequencies convey fine details, allowing the model to effectively disentangle overlapping features. The frequency-domain convolutions allow our model to capture features across multiple projection angles as well, integrating information globally rather than only focusing on local patterns.
Furthermore, we propose a custom loss function that leverages the unique properties of sinograms, including total absorption consistency — reflecting the physical principle that the sum of absorption along each projection line remains constant — and frequency-domain stability, which maintains consistency of structural features across different projection angles, enabling accurate reconstruction of detailed textures and patterns inherent to sinogram data. Specifically, this work consists of three main contributions:
  • A novel approach is designed explicitly for sinogram data by applying frequency domain convolutions, allowing it to disentangle overlapping features inherent to sinograms.
  • A new loss function is developed to align with the unique statistical properties of absorption consistency and frequency consistency of sinograms. This loss function is also closely integrated with the frequency domain convolution, ensuring that the features extracted from the frequency domain, are accurately reflected and preserved.
  • An extensive evaluation is conducted on a real-world dataset [7], and two common synthetic and simulated datasets, demonstrating that FCDM achieves more than 0.95 in SSIM [45] metric and over 30 dB in PSNR metric, with an improvement of up to 33% in SSIM and 29% in PSNR. The subsequent ablation studies then prove the efficacy of our new designs - frequency domain convolutions, and new loss function.
2 Related Work
Deep learning-based methods have shown great promise in RGB image inpainting. [48] proposes a joint optimization method based on a Bayesian framework. [19] applies continuously mask on Transformer-based structure. [9] proposes a new Hourglass attention structure and combines it with the Transformer structure. [21] presents a new inpainting framework Multi-level Interactive Siamese Filtering using an interactive dynamic kernel prediction mechanism. [37] proposes LaMa which utilizes Fast Fourier Convolution [5] to solve large masking problmes. [28] proposes an inpainting method that solely leverages an off-the-shelf unconditionally trained Denoising Diffusion Probabilistic Models (DDPM). [26] uses Diffusion-based structure guidance to solve the problem of semantic discrepancy in masked and unmasked regions. However, when applied to sinograms, these models fall short because they are designed with the assumptions and properties of RGB data, neglecting the feature overlapping problem of sinogram data.
Traditional mathematical methods have also been employed for sinogram inpainting. Linear interpolation [14] and spline interpolation [38] estimate the missing sinogram data based on the known data. Total Variation (TV) Minimization [3], originally developed for image denoising and reconstruction, has also been applied to sinogram inpainting. It works by promoting sparsity in the gradient domain, making it effective for inpainting sinograms with small missing regions. These methods are computationally efficient and easy to implement but overlook essential physical properties and often produce suboptimal results when dealing with large missing areas.
Recently, several deep learning-based models have been proposed for sinogram inpainting. [47, 51] utilize U-Net to directly interpolate missing sinogram data. [39, 23, 40, 46] employ GANs to generate realistic sinogram patches. [27] uses Transformer blocks to inpaint geometric shapes sinograms. These models typically treat the sinograms as normal RGB images without utilizing the specific characteristics of sinogram data. Other deep learning-based models apply unique characteristics of sinograms to fit in sinogram inpainting. [22, 24] incorporate periodicity into the loss function to leverage the cyclical nature of certain sinograms. [44, 23] compare the reconstructed images from the inpainted and ground truth sinograms in their loss calculations. [36] combines coordinate mapping and Fourier feature mapping together with Multilayer Perceptron to help the inpainting of sinogram. However, these models do not utilize efficient feature extraction methods that fit the physical principles of the sinogram generation process.
3 FCDM Design and Analysis
(a) FCDM Overview: In stage 1, the Enc-Dec network is trained to generate a latent representation capturing both spatial and frequency features. In stage 2, with fixed parameters from stage 1, random masking information is embedded using the K, Q, V formulation.
(b) Freq Conv Overview: After obtaining the latent representation from Enc, the results of both standard conv and frequency domain conv applied to the latent features are then concatenated. Finally, a subsequent standard convolution is performed.
Figure 2: Overview of FCDM
FCDM is a two-stage process model that first employs an Encoder-Decoder network to capture both spatial and frequency domain features to better represent the full image by a latent tensor, followed by a Diffusion process that takes mask information and then does inpainting. Fig. 2 shows a brief overview. In stage 1, the Encoder encodes the input sinogram into a latent representation, which is then enhanced using frequency domain convolutions. This enriched latent representation is subsequently decoded to recover the sinogram. Once the training for this stage is done, the Encoder-Decoder network’s parameters are fixed and do not affect the training process for stage 2. Stage 2 involves training a Diffusion Model on the latent space, where various random masks are applied to simulate real-world conditions with missing data. This allows the model to handle complex sinogram inpainting scenarios effectively.
3.1 Frequency Domain Convolution
FCDM begins with the input sinogram being encoded into a latent representation, which is then enhanced through frequency domain operations, and finally decoded back into the reconstructed sinogram.
The Encoder $\mathcal{E}(\cdot)$ transforms the input sinogram $s\in{\mathbb{R}}^{C\times H\times W}$ into a latent representation $h\in{\mathbb{R}}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$:
$h^{C^{\prime}\times H^{\prime}\times W^{\prime}}=\mathcal{E}(s^{C\times H\times W}),$ (1)
where $C$, $H$, and $W$ represent the number of channels, height, and width, respectively. Once the latent representation $h\in{\mathbb{R}}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$ is obtained, we perform a Real Fourier Transform (RFFT) [6] along the width dimension on the latent representation to convert it into the frequency domain. The output $\hat{h}_{w}$ is a complex-valued tensor:
$\hat{h}_{w}=\mathcal{F}_{w}(h)=\Re(\hat{h}_{w})+j\Im(\hat{h}_{w}).$ (2)
Next, we reshape the complex tensor $\hat{h}_{w}$ into a real-valued tensor by concatenating the real and imaginary components along the channel dimension:
${\hat{h}_{w}}^{C^{\prime}\times H^{\prime}\times W^{\prime}}\longrightarrow{\tilde{h}_{w}}^{2C^{\prime}\times H^{\prime}\times\frac{W^{\prime}}{2}}.$ (3)
This operation doubles the number of channels, allowing the model to perform convolution on both components.
In the spatial domain, convolution operations are computationally expensive due to the need to sum over sliding windows of input data. However, by transforming the convolution operation into the frequency domain, we can exploit the Convolution Theorem, which states that convolution in the spatial domain corresponds to Hadamard multiplication in the frequency domain:
$\hat{y}_{w}=\mathcal{G}(\mathcal{K}_{w}\cdot{\tilde{h}_{w}})$ (4)
$\mathcal{G}$ is a non-linear activation function, $\cdot$ denotes Hadamard multiplication, and $\mathcal{K}_{w}$ is the Fourier-transformed convolution kernel. This Hadamard multiplication is computationally more efficient than traditional spatial domain convolution because it avoids the need for sliding windows and reduces the operation to simple multiplications.
After conv, the result is reshaped back to a complex-valued tensor by reconstituting real and imaginary parts:
$\hat{y}_{w}=\hat{y}_{w,\Re}+j\hat{y}_{w,\Im}$ (5)
The complex tensor $\hat{y}_{w}$ is then transformed back into the spatial domain using Inverse Real Fourier Transform [6]:
$y_{w}=\mathcal{F}_{w}^{-1}(\hat{y}_{w})$ (6)
It yields the spatial domain representation $y_{w}$, and similarly, a process is applied along the height dimension and gets $y_{h}$.
The outputs from the frequency domain convolutions along the width and height dimensions are combined:
$y={\alpha}_{w}y_{w}+{\alpha}_{h}\beta y_{h}$ (7)
The contribution from $y_{w}$ emphasizes the frequency characteristics of the sinogram across different angles, capturing essential features within each projection. Conversely, $y_{h}$ focuses on the relationships between different angles, encapsulating the global structure and dependencies across the entire sinogram. The coefficients ${\alpha}_{w}$ and ${\alpha}_{h}$ are selected to balance these two components, ensuring that the model captures both localized frequency information and the overarching global context.
Once the enhanced latent representation $y$ is obtained, it is passed through the Decoder $\mathcal{D}(\cdot)$:
$\overline{s}^{C\times H\times W}=\mathcal{D}(h^{C^{\prime}\times H^{\prime}\times W^{\prime}})$ (8)
This decoding process translates the frequency-enhanced latent features back into the spatial domain, yielding the final inpainted sinogram $\overline{s}\in{\mathbb{R}}^{C\times H\times W}$.
The encoder and decoder remain fixed during the second stage of training, which not only reduces the computational burden but also facilitates easy adaptation for different downstream tasks, such as full CT reconstruction. This design choice enhances the model’s versatility and efficiency in diverse applications.
3.2 Comparison Related to Convolution Locations
There are two strategies for applying frequency domain convolutions in sinogram inpainting: i) performing convolution after downsampling the original sinogram and ii) performing frequency domain convolution within the latent space, post-encoding. The second strategy processes the full-resolution input sinogram through the Encoder, which allows for capturing and preserving a richer and more comprehensive set of features since the encoder captures global features better than downsampling. It enables the model to better capture detailed features for each projection angle as well as the intricate relationships between various angles, resulting in better overall inpainting quality than the first strategy. Besides, performing these convolutions in the latent space rather than on the original images can lower the potential of increase in memory requirements associated with frequency domain convolutions, such as handling complex numbers and full images. More details of ablation evaluation results are shown in Sec. 4.3. Analysis of the complexity of the two strategies is in the supplementary materials.
3.3 Tailored Loss Functions
In designing the loss function for our model, we integrate various established techniques to optimize both perceptual quality and structural accuracy. Pixel loss measures pixel-wise differences to ensure accurate reconstruction. Perceptual loss [49] focuses on maintaining perceptual similarity by comparing high-level features. Adversarial loss [12, 25] uses a discriminator to encourage the generation of realistic images. Codebook loss [41] prevents collapse in the VQ-VAE’s latent codebook by maintaining diversity. Adaptive GAN loss [18] adjusts the adversarial loss weight dynamically, stabilizing the training process.
While these loss functions are highly effective for RGB image inpainting, their direct application to sinogram data is less successful. RGB images are spatially structured and benefit from perceptual and pixel-level losses that align well with human visual perception. In contrast, sinograms represent projection data, which encodes information across different angles and requires consideration of the underlying physical properties [17].
3.3.1 Total Projection Consistency Loss
In the context of CT, one essential physical property is the relationship between the sum of projection data across all angles and the total absorption of the object being imaged. The frequency domain convolution extracts frequency features that encapsulate the physical characteristics of the object, which are inherently linked to its absorption behavior. We first tailor the loss function to enforce this consistency with the total absorption characteristic inherent in CT imaging. By incorporating this term, we ensure that the model not only captures the detailed frequency features but also maintains accurate physical consistency in the inpainted sinogram and further reconstructed images.
The sum of the projection data across all angles ideally equals the total absorption of the object, i.e.:
[…]
where $P_{\theta}(s)$ represents the projection data at angle $\theta$ and position $s$ on the detector, and $\mu(x,y)$ is the linear attenuation coefficient at the point $(x,y)$ in the object.
To enforce this consistency within the loss function, we introduce the total projection consistency loss $\mathcal{L}_{\text{absorp\_sum}}$:
[…]
This loss term ensures that the reconstructed image maintains the physical principle where the total detected radiation equals the total absorption of the object.
The second component of $\mathcal{L}_{\text{absorp\_sum}}$ involves calculating the total absorption of the object using the inverse Radon transform (iRadon transform) [32],
$\mu(x,y)=\mathcal{R}^{-1}\{P_{\theta}(s)\}.$ (11)
3.3.2 Frequency Domain Consistency Loss
The frequency domain loss term directly evaluates the accuracy of the frequency information extracted by the frequency domain convolutional layers. By minimizing the difference between the frequency representations of the inpainted and original sinograms, this loss ensures that the frequency domain convolution effectively captures and preserves the essential frequency characteristics.
To calculate the difference in the frequency domain, we first need to transform the sinograms from the spatial domain to the frequency domain. This transformation is accomplished using the 2D Fourier Transform [6]. For the inpainted sinogram $\hat{\mu}(u,v)$ and the ground truth $\mu(u,v)$, their frequency domain representations $\hat{F}(u,v)$ and $F(u,v)$ are given by:
$\overline{F}(u,v)=\mathcal{F}\{\overline{\mu}(x,y)\}.$ (12)
$F(u,v)=\mathcal{F}\{{\mu}(x,y)\}$ (13)
To quantify this difference and incorporate it into the loss function, we calculate the sum of L2 norm in the frequency domain:
$\mathcal{L}_{\text{freq}}=\sum_{u}\sum_{v}\left\{\overline{F}(u,v)-F(u,v)\right\}^{2}.$ (14)
The final loss function is a weighted combination of several terms, each carefully designed to enforce different aspects of the inpainting process. By integrating multiple loss components, the loss can better leverage the unique characteristics of sinograms, which makes the loss more fit for sinogram inpainting and effectively handles the complex nature of sinogram data.
3.4 Mask Generation and Training
In the second stage of our method, we introduce a masking strategy within the Diffusion process to enhance the model’s robustness and ability to handle complex inpainting conditions. Although actual masking occurs at fixed intervals, these specific intervals can vary in practice. Therefore, FCDM uses random masks in order to adapt to these fluctuations and strengthen the model’s performance in diverse scenarios. Specifically, we automatically generate masks of varying sizes and locations, randomly masking different angles of the sinogram, which simulates the real-world scenario where certain parts of the sinogram data may be missing or corrupted.
To integrate masks into the Diffusion process, we use a Transformer-based mechanism [43], embedding mask information through the Key (K), Query (Q), and Value (V) formulation of the attention mechanism. The Queries are derived from input data, while the Keys and Values come from the mask representation, enabling the model to compute attention scores that highlight relevant features. This adaptive focus helps the model infer missing information effectively, particularly in scenarios with significant missing data, by dynamically guiding the inference process based on input conditions.
4 Evaluation
This section evaluates FCDM on three datasets: Real-world [7], Shapes generated using [42], and Shepp2d generated using [13], and compares it with nine baselines, in which UsiNet [47], SionTx [27] and CoIL [36] are recently works designed for sinogram inpainting, and StrDiffusion [26], CoPaint [48], CMT [19], hourglass [9], RePaint [28], MISF [21], and LaMa [37] are recent works for more general image inpainting. This evaluation has three main objectives: 1) demonstrating FCDM outperforms all baselines, specifically achieving high accuracy on the complex real-world dataset (Sec. 4.2); 2) studying the underlying reasons for performance gains with an ablation study, proving the efficacy of our two key designs (Sec. 4.3); and 3) studying FCDM’s performance with a key factor in practice, mask ratio, showing its robustness (Sec. 4.4).
4.1 Experimental Setup
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.961 (0.917) 0.965 (0.922) 0.954 (0.918) 32.3 (29.5) 31.4 (28.8) 31.4 (28.8)
UsiNet 0.460 (0.416) 0.482 (0.431) 0.494 (0.448) 12.8 (10.7) 14.5 (11.5) 14.9 (12.6)
SionTx 0.572 (0.556) 0.597 (0.544) 0.643 (0.609) 20.4 (18.2) 21.7 (19.5) 21.9 (20.2)
StrDiffusion 0.462 (0.432) 0.508 (0.465) 0.507 (0.474) 13.7 (10.8) 15.7 (14.2) 16.2 (12.4)
CoPaint 0.725 (0.701) 0.738 (0.712) 0.761 (0.717) 23.0 (22.5) 23.2 (21.8) 23.9 (23.2)
CMT 0.436 (0.392) 0.440 (0.407) 0.453 (0.406) 16.1 (14.8) 16.2 (13.2) 16.1 (14.9)
hourglass 0.592 (0.557) 0.601 (0.550) 0.628 (0.581) 21.3 (19.1) 21.7 (19.4) 22.7 (19.2)
RePaint 0.720 (0.682) 0.754 (0.727) 0.766 (0.723) 24.6 (22.3) 25.2 (21.8) 26.3 (23.6)
MISF 0.706 (0.671) 0.717 (0.664) 0.723 (0.699) 25.1 (23.3) 25.2 (23.5) 26.4 (25.1)
LaMa 0.706 (0.672) 0.716 (0.683) 0.715 (0.666) 22.1 (20.6) 22.9 (20.3) 24.6 (22.1)
CoIL 0.555 (0.481) 0.560 (0.544) 0.626 (0.541) 24.7 (22.0) 25.1 (23.0) 26.1 (23.9)
Table 1: Quantitative Accuracy Results: FCDM is compared with all 9 baselines on the Real-world dataset. The values in parentheses correspond to the results of the reconstructed images. The comparisons of Shape and Shepp2d datasets are in supplementary materials.
4.1.1 Implementation Details
We train our model using the Polaris supercomputer at Argonne Leadership Computing Facility (ALCF) resources at Argonne National Laboratory (ANL) [^1] Polaris Cluster at ALCF: https://www.alcf.anl.gov/polaris. Polaris supercomputer consists of 560 compute nodes, each of which is equipped with 4 NVIDIA A100 GPUs (connected via NVLink). We use PyTorch version 2.1.0 and CUDA version 12.2. The dimensions of the input images are 512x512. For frequency domain convolutions, the weights for the $H$ and $W$ dimensions are adjusted to 0.55 and 0.45, respectively. The loss function is a weighted combination of several components with the following coefficient: pixel loss ${\alpha}_{pixel}=1.0$, perceptual loss ${\alpha}_{perceptual}=0.5$, adversarial loss ${\alpha}_{adv}=0.05$, codebook loss ${\alpha}_{cb}=0.5$, absorption consistency loss ${\alpha}_{absorp\_sum}=0.1$, and frequency domain loss ${\alpha}_{freq}=0.1$. During training, the mask ratio is randomly chosen within a range from $0.1$ to $0.9$ with the corresponding proportion of angles being randomly masked. The learning rate is adjusted based on the formula: $lr={lr}_{base}\times batch\,size\times GPU\,num\times grad\,accum\,steps$.
4.1.2 Dataset
Three datasets are used in our experiments, each containing 100k samples with a resolution of $512\times 512$:
Real-world dataset [7] consists of real sinogram images obtained from various materials and objects, derived from actual synchrotron radiation CT scans, including Advanced Photon Source (APS [^2] Advanced Photon Source: https://www.aps.anl.gov) at ANL. This dataset, provided by the x-ray tomography data bank, includes a range of noise levels and artifacts typical in real-world imaging, making it significantly more complex than simulated and synthetic datasets. This real-world dataset is particularly challenging due to the inclusion of dynamic [29] and in situ [31] systems, which are representative of materials imaged at APS and other synchrotron facilities. This complexity offers a rigorous test of the model’s robustness and ability to handle practical challenges in sinogram inpainting.
Shape dataset generated using the scikit-image library [42] contains a variety of simulated geometric shapes, including circles, rectangles, and triangles. This dataset introduces a range of geometric complexities, allowing us to assess the model’s ability to reconstruct shapes with sharp edges and varying curvature.
Shepp2d dataset generated by the Tomopy library [13] comprises simulated shepp-logan phantoms, offering a controlled environment for evaluating the accuracy and stability of reconstruction algorithms. These phantoms are less complex and ideal for testing basic inpainting capabilities.
4.1.3 Metrics
This evaluation leverages two metrics, Structural Similarity Index Measure (SSIM) [45] and Peak Signal-to-Noise Ratio (PSNR), to quantitatively assess the quality of the inpainted sinograms. SSIM assesses structural similarity by considering changes in structure, luminance, and contrast, making it effective for evaluating how well the inpainted sinogram preserves the structural integrity necessary for sinogram inpainting. PSNR measures the ratio between the maximum possible signal power and the noise affecting image fidelity, providing a quantitative assessment of how closely the inpainted sinogram approximates the ground truth, which is crucial for minimizing artifacts.
4.2 Accuracy Comparisons with Baselines
(a) Mask
(b) Ground Truth
(c) Ours
(d) UsiNet
(e) SinoTx
(f) StrDiffusion
(g) Copaint
Figure 3: Visual Results: Lines 1, 3, 5 show visual results of sinogram inpainting of the Real-world dataset, and lines 2, 4, 6 show the reconstructed images of the sinograms. Details of visual results for Shapes and Shepp2d datasets are in supplementary materials.
This section evaluates FCDM’s performance by comparing it against 9 baselines on 3 datasets. For each comparison, we focus on evaluating the SSIM and PSNR metrics specifically within the masked regions, rather than across the entire image. This approach allows us to precisely measure the model’s ability to inpaint the most challenging portions of the sinogram, where data is missing or corrupted. Besides, since the final goal of CT scans is reconstructed images, we also evaluate the images reconstructed by Gridrec[10] on the SSIM and PSNR metrics.
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.961 0.965 0.954 32.3 31.4 31.4
W/o Freq Conv 0.789 0.816 0.813 26.7 27.0 27.3
W/o $H$ Conv 0.846 0.847 0.846 26.1 26.5 27.5
W/o $W$ Conv 0.880 0.884 0.891 27.1 27.9 28.2
Table 2: Ablation Results Related to Freq Conv: The experimental results show that the model’s performance is greatly impacted by the frequency domain convolution. Furthermore, there is a marginally more noticeable drop in performance when the $H$ dimension convolution is absent than when the $W$ dimension convolution is absent. $W$ dimension convolution retrieves important frequency information inside each angle, whereas the $H$ dimension frequency convolution captures the inter-angle correlations.
Tab. 1 summarizes the performance of our model and the baseline models on Real-world dataset [^3] Due to space limitation, please find our results on Shape and Shepp2d datasets in our supplementary materials.. In addition to the quantitative analysis, we also provide visual comparisons between the outputs of our model and the baseline models. Fig. 3 illustrates the inpainted sinograms and reconstructed images using Gridrec [10] when the mask ratio is equal to $0.8$, focusing on showing the visual quality and structural fidelity of the inpainted sinograms and reconstructed images. Our experimental results show that our model significantly outperforms all baselines on the more complex Real-world dataset. The simplicity of the Shapes and Shepp2d datasets allows previous baselines to achieve good results, making our model’s improvements relatively modest (but still obvious) in these cases. Real-world dataset presents greater challenges, and our model excels by effectively leveraging the unique properties of sinogram, leading to a clear and substantial advantage in inpainting quality over the other methods.
4.3 Ablation Studies
Our ablation studies are divided into two key sections, each designed to validate the contributions of a component of our model, frequency domain convolution, and tailored loss functions. More specifically, to assess the impact of frequency domain convolutions, we conducted experiments where this component was either disabled or modified.
4.3.1 Impact of Frequency Domain Convolutions
First, we compare the performance of the model with and without frequency domain convolutions and without $H$ or $W$ dimension convolutions. Results in Tab. 2 clearly show that including frequency domain convolutions significantly enhances the model’s ability to inpaint sinograms with greater accuracy, particularly in challenging scenarios with a significant amount of missing data. More details of visual results are in supplementary materials.
Additionally, we evaluate the locations of the frequency domain convolutions within the model, comparing their application before and after the Encoder. This comparison focuses on the Encoder-Decoder component. Its results in Tab. 3 indicate that applying frequency domain convolutions after the Encoder, within the latent space, leads to better inpainting quality. This is likely due to the better abstraction and compression ability of Encoder compared to downsampling, which allows the frequency domain operations to better capture features from various angles.
To further assess the impact of the use of frequency domain convolution on memory utilization, we also compare peak memory allocation between FCDM incorporating frequency domain convolutions and without it. Our findings reveal no significant difference in memory usage between the two configurations (less than $1\%$ as shown in Supplementary Materials), indicating that the latent space implementation effectively balances memory efficiency and model performance.
Locations SSIM PSNR
Ds Followed by Freq Conv 0.962 33.9
Latent Freq Conv 0.987 35.9
Table 3: Ablation Results Related to Freq Conv Locations: Experiments are done in the first stage. Results show that by conducting frequency domain convolutions on the latent after the input has been passed through, an Encoder produces better performances.
4.3.2 Contribution of Tailored Loss Functions
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.961 0.965 0.954 32.3 31.4 31.4
W/o $L_{absorp\_sum}$ 0.917 0.930 0.927 29.0 29.4 29.6
W/o $L_{freq}$ 0.923 0.935 0.926 29.6 29.4 29.8
Table 4: Ablation Results Related to Tailored Loss: The results illustrate the model’s performance on the real-world dataset when specific terms are removed from the loss function. It is evident that the model’s performance degrades when either $L_{absorp\_sum}$ or $L_{freq}$ is excluded.
The second part of our ablation study examines the role of the tailored loss functions, specifically the absorption consistency loss and the frequency domain loss. We conduct experiments where each of these losses is individually removed from the training process. Results in Tab. 4 demonstrate that the tailored losses are essential for achieving high-quality reconstructions. Removing the absorption consistency loss results in a failure to fully adhere to the physical properties of CT imaging. Similarly, the absence of the frequency domain loss results in a degradation in the ability to maintain the correct frequency characteristics, leading to less accurate inpainting.
4.4 Mask Ratio Analysis
Figure 4: Relationships between Mask Ratio and Metrics: As mask ratios increase, SSIM and PSNR tend to rise but are not strictly monotonic, exhibiting fluctuations.
During our experiments, we observe an intriguing trend: as the mask ratio increases, the SSIM and PSNR metrics in the masked regions of the inpainted sinograms also tend to improve. Fig. 4 shows the plot of this trend of the real-world dataset. This phenomenon can be attributed to the model’s ability to better capture and learn global features when faced with larger masked areas. With a higher mask ratio, the model is less likely to overfit to local features.
5 Conclusion and Future Work
This work introduces a new model FCDM for sinogram inpainting that applies frequency domain convolutions and tailored loss functions. The evaluation results demonstrate that FCDM outperforms state-of-the-art existing methods, achieving SSIM over 0.95 and PSNR above 30 dB (up to $33\%$ gains in SSIM and $29\%$ in PSNR). The quality of the reconstructed images based on the output of FCDM is close to the ground truth. The following ablation studies further validate the efficacy of each designed component, frequency domain processing, and specialized loss functions. This study provides a foundation for extending our approach to complete CT reconstruction. In future work, we aim to develop the model further to take sinograms as input and generate fully reconstructed CT images directly.
Supplementary Material
6 Complexity Analysis of Different Convolution Locations
6.1 Downsampling Followed by Frequency Domain Convolution
Downsampling reduces the spatial dimensions of the input image from $H\times W$ to $\frac{H}{r}\times\frac{W}{r}$, where $r$ is the downsampling factor:
$\mathcal{O}_{down}=\mathcal{O}\left(\frac{2CWH}{r^{2}}\right).$ (15)
After downsampling, half of the channels undergoes standard convolution, with complexity:
$\mathcal{O}_{s}=\mathcal{O}\left(C^{2}k_{s}^{2}\frac{WH}{r^{2}}\right),$ (16)
where $k_{s}$ represents the size of the convolution kernel.
For the other channels, a Fourier Transform is applied, followed by a frequency domain convolution, which is equal to Hadamard multiplication in the spatial domain, and then an inverse Fourier Transform:
$\begin{split}\mathcal{O}_{f}&=\mathcal{O}\left(C\frac{WH}{r^{2}}\log{\frac{WH}{r^{2}}}\right)+\mathcal{O}\left(C\frac{WH}{r^{2}}\right).\\ \end{split}$ (17)
Then a convolution is performed to restore the channel count:
$\mathcal{O}_{r}=\mathcal{O}\left(2C^{2}k_{r}^{2}\frac{WH}{r^{2}}\right)$ (18)
Finally, the result after the restore convolution is passed through an Encoder $\mathcal{E}(\cdot)$. The complexity of it depends on the number of layers $L$, the number of channels $C_{i}$ in each layer, and the size of the convolutional kernels $k_{i}$:
$\mathcal{O}_{e}=\mathcal{O}\left(\sum_{i=1}^{L}C_{i-1}C_{i}k_{i}^{2}\frac{H_{i}W_{i}}{r^{2}}\right),$ (19)
where $H_{i}$ and $W_{i}$ are the dimensions of the feature map at that layer.
6.2 Latent Space Frequency Domain Convolution
The input image is first passed into an Encoder $\mathcal{E}(\cdot)$, which transforms it into a latent representation $h$ with dimensions $C_{L}\times H_{L}\times W_{L}$:
$\mathcal{O}_{e}=\mathcal{O}\left(\sum_{i=1}^{L}C_{i-1}C_{i}k_{i}^{2}H_{i}W_{i}\right).$ (20)
After encoding, standard convolution is applied to the latent:
$\mathcal{O}_{s}=\mathcal{O}(C_{L}^{2}k_{s}^{2}H_{L}W_{L}).$ (21)
Frequency domain convolution followed by an inverse Fourier Transform are also applied:
$\begin{split}\mathcal{O}_{f}&=\mathcal{O}(C_{L}H_{L}W_{L}\log{H_{L}W_{L}})+\mathcal{O}(C_{L}H_{L}W_{L}).\\ \end{split}$ (22)
Finally, a convolution to restore the channel count is performed:
$\mathcal{O}_{r}=\mathcal{O}(2C_{L}Ck_{r}^{2}H_{L}W_{L})$ (23)
7 Supplementary Evaluation Results
(a) Ours
(b) W/o Freq Conv
Figure 5: Visual Ablation Results: Results without frequency domain convolution are blurred.
Methods Peak Memory Usage (MB)
Ours 21497.01
W/o Freq Conv 21464.02
Table 5: Peak Memory Usage Comparison: The difference in memory usage between the two configurations is less than $1\%$, indicating that the implementation of frequency-domain convolutions in the latent space is memory efficient
(a) Mask
(b) Ground Truth
(c) Ours
(d) UsiNet
(e) SinoTx
(f) StrDiffusion
(g) CoPaint
Figure 6: Visual Results of Shape and Shepp2d Datasets: Line 1 and 3 are visual results of sinogram inpainting of Shape dataset and Shepp2d datset, respectively. Line 2 and 4 are visual results of the reconstructed images of the sinograms using iRadon[32]. The visual comparisons highlight our model’s ability to preserve structural details and reduce artifacts.
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.965 (0.923) 0.977 (0.937) 0.977 (0.954) 36.2 (34.7) 36.3 (35.2) 36.6 (34.2)
UsiNet 0.509 (0.480) 0.549 (0.511) 0.546 (0.512) 14.1 (10.9) 13.7 (12.6) 15.3 (12.8)
SionTx 0.746 (0.717) 0.743 (0.696) 0.796 (0.750) 26.5 (25.1) 26.9 (25.2) 27.1 (25.1)
StrDiffusion 0.524 (0.497) 0.552 (0.506) 0.600 (0.572) 14.3 (12.1) 15.1 (13.5) 15.9 (13.5)
CoPaint 0.903 (0.868) 0.929 (0.889) 0.908 (0.888) 30.2 (28.6) 31.0 (29.3) 33.0 (31.3)
CMT 0.454 (0.413) 0.469 (0.427) 0.477 (0.431) 19.9 (17.1) 20.2 (17.4) 21.0 (18.1)
hourglass 0.807 (0.772) 0.813 (0.781) 0.823 (0.781) 27.1 (24.7) 27.5 (24.3) 28.2 (26.5)
RePaint 0.915 (0.869) 0.918 (0.873) 0.926 (0.899) 33.1 (32.3) 33.6 (31.7) 33.69 (31.2)
MISF 0.716 (0.692) 0.739 (0.717) 0.753 (0.735) 26.4 (25.3) 27.6 (26.5) 27.8 (25.0)
LaMa 0.747 (0.718) 0.762 (0.730) 0.779 (0.726) 26.1 (23.9) 27.8 (25.8) 28.6 (27.8)
CoIL 0.746 (0.683) 0.786 (0.709) 0.809 (0.734) 27.9 (25.3) 28.1 (25.7) 29.0 (26.6)
Table 6: Quantitative Accuracy Results on the Shape Dataset: The results demonstrate that our model excels in handling sinograms of objects with sharp edges and geometric structures, outperforming the baselines in the inpainting performance.
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.966 (0.936) 0.978 (0.954) 0.977 (0.961) 37.7 (36.6) 37.2 (34.5) 37.8 (35.5)
UsiNet 0.494 (0.447) 0.499 (0.462) 0.492 (0.452) 13.1 (12.1) 15.0 (13.2) 15.6 (14.5)
SionTx 0.967 (0.938) 0.968 (0.926) 0.962 (0.906) 37.2 (35.8) 37.5 (34.4) 36.3 (33.9)
StrDiffusion 0.516 (0.497) 0.554 (0.505) 0.562 (0.538) 15.5 (12.5) 16.3 (14.9) 16.7 (15.3)
CoPaint 0.932 (0.892) 0.947 (0.891) 0.946 (0.914) 32.0 (30.7) 31.8 (30.5) 32.7 (32.0)
CMT 0.504 (0.457) 0.507 (0.477) 0.518 (0.485) 20.8 (19.3) 22.6 (19.3) 22.7 (19.5)
hourglass 0.826 (0.770) 0.843 (0.813) 0.851 (0.818) 27.5 (24.9) 27.7 (26.3) 28.1 (26.9)
RePaint 0.912 (0.877) 0.931 (0.894) 0.934 (0.913) 33.2 (31.3) 34.3 (31.2) 35.2 (33.5)
MISF 0.731 (0.702) 0.749 (0.713) 0.743 (0.720) 25.8 (24.1) 26.5 (23.7) 26.9 (25.4)
LaMa 0.698 (0.663) 0.708 (0.686) 0.720 (0.692) 25.0 (23.6) 25.8 (23.0) 26.6 (23.6)
CoIL 0.814 (0.749) 0.855 (0.795) 0.870 (0.809) 28.9 (27.8) 29.1 (28.0) 30.1 (28.8)
Table 7: Quantitative Accuracy Results on the Shape2d Dataset: The controlled and simplified nature of the Shepp2d dataset enables previous baselines to perform well, yet our model still demonstrates clear improvements in the inpainting performance.
7.1 Supplementary Comparison Results
Tab. 6Tab. 7, and Fig. 6 present the quantitative and qualitative comparisons on the Shape and Shepp2d datasets against several state-of-the-art baselines. For both datasets, FCDM consistently achieves higher SSIM and PSNR scores compared to previous baselines, indicating superior inpainting accuracy of our model.
On the Shape dataset, which consists of geometric shapes with distinct edges and simple textures, our model demonstrates strong performance in accurately completing sinograms of objects with sharp boundary features. For the Shepp2d dataset, despite its simpler structure, the results confirm that our model can still effectively capture its features and maintain high-quality inpainting. Although the improvements in these controlled datasets are relatively modest, they showcase our model’s robustness and consistency in diverse scenarios, demonstrating that it can perform on par with or better than previous baselines.
Fig. 5 provides visual results from the ablation studies conducted on the Real-world dataset, offering a comprehensive analysis of the contributions of frequency-domain convolutions. The visual comparisons clearly illustrate that removing the frequency-domain convolutions leads to notable degradation in inpainting quality. In the absence of frequency-domain convolutions, the model struggles to capture fine-grained details and often introduces visible artifacts in the reconstructed sinogram, which subsequently affects the quality of the reconstructed images. By contrast, when frequency-domain convolutions are included, the model successfully disentangles mixed-frequency components, preserves structural coherence across projection angles, and can further reduce reconstruction errors. These findings provide strong evidence for the effectiveness of incorporating frequency-domain information, highlighting its critical role in enhancing the model’s ability to address the inherent challenges of sinogram inpainting.
7.2 Memory Usage Caused by Frequency Domain Convolution
We conduct an additional experiment to thoroughly assess the impact of frequency-domain convolutions on memory usage. Specifically, we measure the peak memory allocation for FCDM during the inference phase, on two configurations: with and without frequency-domain convolutions. The experiment is performed using a batch size of 16 samples, each with a resolution of $512\times 512$. The results in Tab. 5 indicate that incorporating frequency-domain convolutions does not lead to a significant increase in memory consumption. This demonstrates that the additional computation is offset by applying the frequency-domain convolutions in the latent space instead of the whole image. Our findings further confirm that the proposed method maintains a balance between accuracy and resource utilization.