### v3

# FCDM: A Physics-Guided Bidirectional Frequency Aware Convolution and Diffusion-Based Model for Sinogram Inpainting

Abstract
Computed tomography (CT) is widely used in industrial and medical imaging, but sparse-view scanning reduces radiation exposure at the cost of incomplete sinograms and challenging reconstruction. Existing RGB-based inpainting models struggle with severe feature entanglement, while sinogram-specific methods often lack explicit physics constraints. We propose FCDM, a physics-guided, frequency-aware sinogram inpainting framework. It integrates bidirectional frequency-domain convolutions to disentangle overlapping features while enforcing total absorption and frequency-domain consistency via a physics-informed loss. To enhance diffusion-based restoration, we introduce a Fourier-enhanced mask embedding to encode angular dependencies and a frequency-adaptive noise scheduling strategy that incorporates a soft row-wise absorption constraint to maintain physical realism. Experiments on synthetic and real-world datasets show that FCDM outperforms existing methods, achieving SSIM over 0.95 and PSNR above 30 dB, with up to 33% and 29% improvements over baselines.
1 Introduction
X-ray computed tomography (XCT) is a vital 3D imaging technique widely used in synchrotron facilities for high-resolution internal structure analysis. For instance, the nano-CT beamlines at the Advanced Photon Source (APS) at Argonne National Laboratory (ANL) offer cutting-edge imaging capabilities, achieving sub-20 nm spatial resolution [7]. These capabilities enable critical studies in battery defect analysis [44, 29], neurodegenerative disease research [19, 12, 16], and materials science [32]. XCT generates parallel beam projections by rotating a sample under collimated X-rays. These projections are organized into a 3D dataset and sliced row-by-row, producing sinograms that are then reconstructed into 3D images via computationally intensive algorithms [4, 15, 1]. While XCT provides high-precision imaging, its high radiation dose can degrade samples, increase scanning time, and raise computational costs [2]. Sparse-view CT mitigates these issues by reducing the number of projections, but this results in limited measurements, leading to severe noise, artifacts and structural loss in reconstructed images. Thus, effective sinogram inpainting is essential for restoring missing projections and improving CT reconstruction quality.
(a) RGB Img
(b) RGB Spec
(c) Sino Img
(d) Sino Spec
Figure 1: Comparisons of RGB and sinogram and their spectra. Unlike RGB images, which have localized frequency components, sinograms exhibit structured spectral distributions due to the Radon transform.
One key limitation of conventional spatial-domain inpainting models [25, 43, 18, 10, 27, 21, 34] is their inability to handle feature entanglement [9] in sinograms. Unlike RGB images, where each pixel corresponds to a localized structure, sinogram pixels encode accumulated attenuation values along X-ray projection paths, leading to feature entanglement. Consequently, missing sinogram regions introduce ambiguities that propagate through reconstruction, making spatial-based inpainting ineffective. This issue is further compounded by the unique spectral characteristics of sinograms. While RGB images have localized frequency components, sinograms exhibit structured spectral distributions due to the Radon transform [31], encoding long-range dependencies across projection angles (Fig. 1). These combined factors make traditional spatial inpainting methods ineffective, underscoring the necessity of frequency-domain processing. To address this, we introduce bidirectional frequency-domain convolutions (BFDC), which independently process frequency components along both the projection angle and detector dimensions. Unlike purely spatial convolutions, which struggle with mixed projection components, and purely frequency-based approaches, which lose spatial locality, BFDC combines the advantages of both domains. It processes spatial features to retain local structure and also applies spectral decomposition to disentangle overlapping features, enabling more effective inpainting. Instead of applying convolutions in a single dimension or fully in 2D, our bidirectional approach independently extracts features along projection angles and detector positions, balancing global and local dependencies. By assigning different weights along each dimension, BFDC dynamically preserves global correlations while ensuring local feature extractions, leading to more accurate sinogram restoration.
Another key limitation of existing methods is their failure to enforce physical consistency, particularly total absorption conservation [17]—–a fundamental X-ray imaging principle requiring the sum of attenuation values along a projection to remain stable. Most deep learning-based inpainting models treat sinograms as generic images, optimizing for pixel fidelity while neglecting physical constraints. To address this, we introduce a physics-guided loss function that preserves total absorption consistency and frequency-domain stability, ensuring that restored sinograms adhere to physical principles and further reduce artifacts.
Further, standard diffusion models use binary 0/1 masks, which are effective for RGB images where missing pixels correspond to localized features. However, in sinograms, missing regions are correlated across angles, as the same structure appears at different locations. A standard 0/1 mask does not indicate the projection angle, making it challenging for the model to infer how to restore structures. By incorporating Fourier-enhanced mask embedding (FEME), our approach enables the model to distinguish missing regions based on their projection angles, allowing it to leverage observed structures from other angles for more consistent and physically plausible restoration.
Finally, standard diffusion models apply uniform Gaussian noise across all frequency components, ignoring that recovery prioritizes global structures before fine details. To align with this process, we introduce frequency-adaptive noise scheduling (FANS), applying stronger noise to high-frequency components early on to disrupt fine details first, while later steps focus on low-frequency noise to preserve structural stability. Additionally, instead of enforcing strict absorption consistency, we introduce a soft constraint to prevent excessive deviations in total absorption. This avoids the model wasting capacity on correcting non-physical intensity shifts, allowing it to focus on meaningful sinogram inpainting. By permitting controlled absorption variations, our approach ensures both physics alignment and generalization to real-world scanning variations.
In summary, our key contributions are:
  • We propose FCDM, a novel physics-guided bidirectional frequency-aware diffusion-based model tailored for sinogram inpainting, addressing both feature disentanglement and physics-compliance challenges.
  • We introduce bidirectional frequency-domain convolutions to enhance feature extraction across projection angles and detector dimensions, and design a physics-informed loss function enforcing total absorption and frequency consistency.
  • We improve the robustness of diffusion-based inpainting by incorporating Fourier-enhanced mask embedding and frequency-adaptive noise scheduling, enabling better generalization across varying sinogram corruption patterns.
  • An extensive evaluation is conducted on a real-world dataset [8], and two common synthetic and simulated datasets, demonstrating that FCDM achieves more than 0.95 in SSIM [40] metric and over 30 dB in PSNR metric, with an improvement of up to 33% in SSIM and 29% in PSNR. The subsequent ablation studies then prove the efficacy of our new designs.
2 Related Work
Deep learning-based methods have shown great promise in RGB image inpainting. [43] proposes a joint optimization method based on a Bayesian framework. [18] applies a continuous mask on Transformer-based structure. [10] proposes a new Hourglass attention structure and combines it with the Transformer structure. [21] presents a new inpainting framework with multi-level interactive siamese filtering using an interactive dynamic kernel prediction mechanism. [34] proposes LaMa which utilizes fast Fourier convolution [5] to solve large masking problmes. [27] proposes an inpainting method that solely leverages an off-the-shelf unconditionally trained Denoising Diffusion Probabilistic Models (DDPM). [25] uses diffusion-based structure guidance to solve the problem of semantic discrepancy in masked and unmasked regions. However, RGB image inpainting methods do not consider the unique physical constraints and feature entanglements in sinograms, making them less effective for sinogram inpainting. Our method is specifically designed for sinogram data, incorporating frequency-domain processing and physics-guided constraints to improve inpainting accuracy.
Traditional mathematical methods have also been employed for sinogram inpainting. Linear interpolation [14] and spline interpolation [35] estimate the missing sinogram data based on the known data. Total Variation (TV) minimization [3], originally developed for image denoising and reconstruction, has also been applied to sinogram inpainting. It works by promoting sparsity in the gradient domain, making it effective for inpainting sinograms with small missing regions. While these mathematical methods provide simple and interpretable solutions, they struggle with large missing regions and fail to leverage the global structure of sinograms. Our approach overcomes these limitations by employing a diffusion-based model with frequency-aware inpainting, enabling more accurate and structurally consistent reconstructions.
Recently, several deep learning-based models have been proposed for sinogram inpainting. [42, 45] utilize U-Net to directly interpolate missing sinogram data. [36, 23, 37, 41] employ GANs to generate realistic sinogram patches. [26] uses Transformer blocks to inpaint geometric shapes sinograms. Unlike these methods, which rely on spatial-domain interpolation, our approach integrates frequency-domain convolutions and a physics-aware loss to ensure physically consistent inpainting.
Other deep learning-based models apply unique characteristics of sinograms. [22, 24] incorporate periodicity into the loss function to leverage the cyclical nature of certain sinograms. [39, 23] compare the reconstructed images from the inpainted and ground truth sinograms in their loss calculations. [33] combines coordinate mapping and Fourier feature mapping together with multilayer perceptron to help the inpainting of sinogram. Our method differs by explicitly modeling sinogram characteristics in the frequency domain while also integrating Fourier-enhanced mask embeddings and frequency-adaptive noise scheduling to enhance inpainting robustness and quality.
3 FCDM Design and Analysis
Figure 2: FCDM Overview: In stage 1, the enc-dec network is trained to generate a latent representation capturing both spatial and frequency features. In stage 2, with fixed parameters from stage 1, random masking information is embedded.
FCDM is a two-stage process model that first employs an encoder-decoder network to capture both spatial and frequency domain features, followed by a diffusion-based process for inpainting using mask information. Fig. 2 shows a brief overview. In stage 1, the encoder encodes the input sinogram into a latent representation, which is then enhanced using BFDC to extract structural features across projection angles and detector positions. This enriched latent representation is subsequently decoded to recover the sinogram. Once this stage is trained, the encoder-decoder network’s parameters are fixed, ensuring stable feature extraction during the next phase. In stage 2, a diffusion-based model is trained on the latent space with random masks applied to simulate real-world missing data conditions. To guide the inpainting process, we incorporate FEME, allowing the model to better distinguish missing regions based on their projection geometry. Additionally, we introduce FANS, where early diffusion steps apply more noise to high-frequency components, while later steps increase noise injection into low-frequency components. This aligns the noise injection process with the natural denoising sequence, improving inpainting stability.
3.1 Bidirectional Frequency-Domain Convolutions
To enhance feature extraction in the latent space, we introduce BFDC. Given a sinogram input $s\in\mathbb{R}^{C\times H\times W}$, the encoder $\mathcal{E}(\cdot)$ maps it to a latent representation $h\in\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$:
$h=\mathcal{E}(s).$ (1)
Next, both spatial and frequency-domain convolutions are applied to refine feature extraction. First, a standard spatial convolution $\mathcal{G}(\cdot)$ is performed:
$h_{s}=\mathcal{G}(h),$ (2)
where $\mathcal{G}(\cdot)$ represents a spatial convolutional operator. We also perform frequency-domain convolutions to complement spatial feature extraction. We apply a real Fourier transform (RFFT) [6] along both the width $W\prime$ and height $H\prime$ dimensions, converting the latent representation into the frequency domain:
$\hat{h}_{w}=\mathcal{F}_{w}(h),\quad\hat{h}_{h}=\mathcal{F}_{h}(h).$ (3)
Both transformations produce complex-valued tensors, which are decomposed into their real and imaginary components before convolution. The frequency-domain convolution is then performed using Hadamard multiplication, computationally more efficient than spatial convolution:
$\hat{y}_{w}=\mathcal{K}_{w}\cdot\hat{h}_{w},\quad\hat{y}_{h}=\mathcal{K}_{h}\cdot\hat{h}_{h}.$ (4)
where $\mathcal{K}_{w}$ and $\mathcal{K}_{h}$ are Fourier-transformed convolution kernels. The processed representations are then mapped back to the spatial domain using an inverse RFFT:
$y_{w}=\mathcal{F}_{w}^{-1}(\hat{y}_{w}),\quad y_{h}=\mathcal{F}_{h}^{-1}(\hat{y}_{h}).$ (5)
Finally, spatial and frequency-domain features are fused, ensuring that both spatial and frequential characteristics are preserved:
$y=\alpha_{s}h_{s}+\alpha_{w}y_{w}+\alpha_{h}y_{h}.$ (6)
The weighting factors $\alpha_{s}$, $\alpha_{w}$ and $\alpha_{h}$ ensure that spatial-domain information and frequency-domain information along both projection angles and detector positions are optimally integrated, allowing the model to effectively disentangle overlapping features while maintaining projection consistency. The decoder $\mathcal{D}(\cdot)$ then reconstructs the inpainted sinogram:
$\overline{s}=\mathcal{D}(y).$ (7)
3.2 Physics-Guided Loss Functions
To ensure physically consistent inpainting, we introduce two loss functions: total projection consistency loss and frequency domain consistency loss.
3.2.1 Total Projection Consistency Loss
In CT, the total X-ray absorption should be preserved across all projection angles. This ensures that the reconstructed sinogram does not introduce non-physical intensity shifts, which could propagate into CT reconstruction. To enforce this, we define the total absorption constraint:
[…]
where $P_{\theta}(s)$ represents the projection data at angle $\theta$ and detector position $s$, while $\mu(x,y)$ denotes the object’s linear attenuation coefficient. We incorporate this constraint into the loss function as:
[…]
where $\mathcal{R}^{-1}$ represents the inverse Radon transform used to reconstruct the attenuation map. This loss encourages the model to maintain global absorption consistency in the inpainted sinogram.
3.2.2 Frequency Domain Consistency Loss
Since sinogram information is naturally structured in the frequency domain, ensuring spectral consistency between inpainted and ground-truth sinograms is essential. We define this constraint by minimizing the L2 norm of the frequency difference between the reconstructed and ground-truth sinograms:
$\mathcal{L}_{\text{freq}}=\sum_{u}\sum_{v}\left\{\mathcal{F}(\overline{\mu}(x,y))-\mathcal{F}({\mu}(x,y))\right\}^{2}.$ (10)
where $\mathcal{F}$ denotes the 2D Fourier transform. This loss ensures that the model accurately captures the structural characteristics of sinograms in the frequency domain, preventing spectral artifacts.
3.2.3 Overall Loss
The final loss for stage 1 training combines pixel-wise reconstruction, physics-guided constraints, and adversarial learning:
$\mathcal{L}_{1}=\lambda_{\text{pixel}}\mathcal{L}_{\text{pixel}}+\lambda_{\text{adv}}\mathcal{L}_{\text{adv}}+\lambda_{\text{absorp}}\mathcal{L}_{\text{absorp}}+\lambda_{\text{freq}}\mathcal{L}_{\text{freq}}.$ (11)
where $\mathcal{L}_{\text{pixel}}$ ensures pixel-wise reconstruction accuracy, and $\mathcal{L}_{\text{adv}}$ is the adversarial loss. In stage 2, the diffusion model optimizes a mean squared error (MSE) loss.
3.3 Fourier-Enhanced Mask Embedding
Standard diffusion-based inpainting methods use binary 0/1 masks to indicate missing regions, which is sufficient for natural images but inadequate for sinograms. Unlike RGB images, where missing pixels can be inferred based on local spatial continuity, sinograms exhibit global feature entanglement due to the projection process. The same physical structure can appear at different positions across projection angles, yet a simple binary mask provides no information about these angular relationships. As a result, the model may treat missing regions as independent spatial holes, failing to recognize that certain missing areas correspond to the same structure observed at different angles.
To address this, we introduce FEME, where the mask is augmented with projection angle information using Fourier encoding. Given a projection angle $\theta$, we compute its Fourier representation as
$\Phi(\theta)=\left[\sin(2\pi f_{1}\theta),\cos(2\pi f_{1}\theta),\dots,\cos(2\pi f_{N}\theta)\right],$ (12)
where $f_{i}$ represents a set of predefined frequencies. The enhanced mask representation is then obtained via element-wise Hadamard multiplication:
$M^{\prime}(\theta)=M(\theta)\odot[1;\Phi(\theta)].$ (13)
This transformation ensures that each missing region retains its original binary mask values while incorporating angular dependencies, making the mask projection-aware. To integrate this enhanced mask representation into the diffusion model, we employ an attention-based mechanism, where the encoded mask features interact with the latent space via key-query-value attention. By embedding this angular information, the model can better recognize correspondences between missing regions across different angles, leading to more coherent and physically consistent restorations in sinogram inpainting.
3.4 Frequency-Adaptive Noise Scheduling
In standard diffusion models, noise is applied uniformly across all frequency components, disregarding the varying importance of different spectral bands in sinogram reconstruction. To enhance the effectiveness of the denoising process, we introduce FANS that dynamically adjusts noise injection across different spectral regions over the course of diffusion.
The process begins by generating standard Gaussian noise $\epsilon_{t}\sim\mathcal{N}(0,I)$, followed by a two-dimensional Fourier transform to obtain the frequency representation of the noise,
$\epsilon_{t}^{\text{freq}}=\mathcal{F}(\epsilon_{t}).$ (14)
Instead of applying uniform noise, we modulate its intensity in the frequency domain by introducing a time-dependent frequency weighting function $m_{t}(u,v)$, which controls the distribution of noise across different frequency components. Specifically, the weighting function is formulated as
$m_{t}(u,v)=1+\alpha\left(1-\frac{t}{T}\right)\cdot\text{dist}(u,v)+\beta\frac{t}{T}\cdot(1-\text{dist}(u,v)),$ (15)
where $\text{dist}(u,v)=\frac{\sqrt{u^{2}+v^{2}}}{d_{\max}}$ represents the normalized distance of the frequency component from the lowest frequency, with
$d_{\max}=\sqrt{(H/2)^{2}+(W/2)^{2}}$ (16)
denoting the maximum frequency distance. The parameters $\alpha$ and $\beta$ control the relative enhancement of high-frequency and low-frequency noise, respectively, while $T$ is the total number of diffusion steps and $t$ denotes the current step.
The weighted noise is obtained by element-wise multiplication in the frequency domain,
$\epsilon_{t}^{\text{weighted}}=m_{t}\cdot\epsilon_{t}^{\text{freq}},$ (17)
followed by an inverse Fourier transform to recover the adjusted noise in the spatial domain,
$\epsilon_{t}^{\text{adapted}}=\mathcal{F}^{-1}(\epsilon_{t}^{\text{weighted}}).$ (18)
While this adaptive noise scheduling improves the quality of the denoising process, it does not inherently preserve the physical consistency of sinograms, particularly in terms of total absorption along projection paths. In standard diffusion models, the added noise may disrupt the total attenuation consistency, leading to inconsistencies that require the model to correct unrealistic variations rather than focus on meaningful reconstruction. To address this, we introduce a soft projection consistency constraint that prevents excessive deviation in total absorption without imposing a strict constraint that could hinder generalization.
To enforce this constraint, we first compute the row-wise sum of the adapted noise, given by
[…]
where $i$ indexes the sinogram rows. To mitigate large deviations, we adjust the noise using a row-wise normalization approach,
$\epsilon_{t}^{\prime}=\epsilon_{t}^{\text{adapted}}-\frac{\text{noise\_sum}}{W}\cdot\mathbf{1},$ (20)
where $W$ is the width of the sinogram and $\mathbf{1}$ represents a unit vector to ensure uniform correction across each row.
To avoid an overly rigid constraint, we introduce a soft weighting factor $\lambda$, allowing a controlled level of deviation:
$\epsilon_{t}^{\text{final}}=\epsilon_{t}^{\text{adapted}}-\lambda\frac{\text{noise\_sum}}{W}\cdot\mathbf{1}.$ (21)
When $\lambda=1$, the total absorption is fully preserved, while $\lambda=0$ ignores consistency entirely. Selecting $10<\lambda<$ 1 provides a balance, enabling the model to maintain reasonable physical properties while retaining flexibility in handling real-world variations.
Finally, the frequency-adaptive noise, now adjusted for projection consistency, is injected into the diffusion process following the standard formulation,
$x_{t}=\sqrt{1-\beta_{t}}x_{t-1}+\sqrt{\beta_{t}}\epsilon_{t}^{\text{final}}.$ (22)
4 Evaluation
This section evaluates FCDM on three datasets: Real-world [8], Shapes generated using [38], and Shepp2d generated using [13], and compares it with nine baselines. Among these, UsiNet [42], SionTx [26], and CoIL [33] are recent methods designed for sinogram inpainting, while StrDiffusion [25], CoPaint [43], CMT [18], Hourglass [10], RePaint [27], MISF [21], and LaMa [34] are general image inpainting approaches.
The evaluation has three main objectives: (1) demonstrating that FCDM outperforms all baselines, particularly on complex real-world datasets (Sec. 4.2); (2) conducting ablation studies to validate the effectiveness of our four key designs, including bidirectional frequency-domain convolutions, physics-guided loss function, Fourier-enhanced mask embedding, and frequency-adaptive noise scheduling (Sec. 4.3).
4.1 Experimental Setup
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.959 (0.916) 0.964 (0.921) 0.953 (0.919) 32.1 (29.4) 31.5 (28.9) 31.3 (28.7)
UsiNet 0.460 (0.416) 0.482 (0.431) 0.494 (0.448) 12.8 (10.7) 14.5 (11.5) 14.9 (12.6)
SionTx 0.572 (0.556) 0.597 (0.544) 0.643 (0.609) 20.4 (18.2) 21.7 (19.5) 21.9 (20.2)
StrDiffusion 0.462 (0.432) 0.508 (0.465) 0.507 (0.474) 13.7 (10.8) 15.7 (14.2) 16.2 (12.4)
CoPaint 0.725 (0.701) 0.738 (0.712) 0.761 (0.717) 23.0 (22.5) 23.2 (21.8) 23.9 (23.2)
CMT 0.436 (0.392) 0.440 (0.407) 0.453 (0.406) 16.1 (14.8) 16.2 (13.2) 16.1 (14.9)
hourglass 0.592 (0.557) 0.601 (0.550) 0.628 (0.581) 21.3 (19.1) 21.7 (19.4) 22.7 (19.2)
RePaint 0.802 (0.738) 0.811 (0.769) 0.805 (0.762) 25.5 (23.3) 25.8 (22.7) 26.1 (24.0)
MISF 0.790 (0.730) 0.801 (0.722) 0.798 (0.748) 25.7 (24.2) 25.9 (24.4) 26.2 (25.0)
LaMa 0.779 (0.732) 0.791 (0.745) 0.785 (0.723) 23.8 (22.0) 24.1 (21.8) 25.0 (23.5)
CoIL 0.555 (0.481) 0.560 (0.544) 0.626 (0.541) 24.7 (22.0) 25.1 (23.0) 26.1 (23.9)
Table 1: Quantitative Accuracy Results: FCDM is compared with all 9 baselines on the Real-world dataset. The values in parentheses correspond to the results of the reconstructed images. The comparisons of Shape and Shepp2d datasets are in supplementary materials.
4.1.1 Implementation Details
We train our model using the Polaris supercomputer at Argonne Leadership Computing Facility (ALCF) resources at Argonne National Laboratory (ANL) [^1] Polaris Cluster at ALCF: https://www.alcf.anl.gov/polaris. Polaris supercomputer consists of 560 compute nodes, each of which is equipped with 4 NVIDIA A100 GPUs (connected via NVLink). We use PyTorch version 2.1.0 and CUDA version 12.2. The dimensions of the input images are $512\times 512$. For frequency domain convolutions, the weights for the $H$ and $W$ dimensions are adjusted to 0.55 and 0.45, respectively, as sinograms exhibit smoother variations along the projection angle direction ($H$) but more complex frequency distributions along the detector axis ($W$). Assigning slightly higher weight to $H$ ensures stable training and better global feature extraction while preventing overfitting to local high-frequency noise in $W$ [20]. The loss function is a weighted combination of several components with the following coefficient: pixel loss ${\alpha}_{pixel}=1.0$, adversarial loss ${\alpha}_{adv}=0.5$, absorption consistency loss ${\alpha}_{absorp\_sum}=0.1$, and frequency domain loss ${\alpha}_{freq}=0.1$. For FEME, we set the Fourier encoding frequencies as $f_{i}=i$ with $N=4$. For FANS, we use $\alpha=1.2$ for high-frequency perturbation scaling and $\beta=0.6$ for low-frequency scaling. The soft absorption constraint is controlled by $\lambda=0.3$, allowing limited variations in total absorption to improve generalization. During training, the mask ratio is randomly chosen within a range from $0.1$ to $0.9$ with the corresponding proportion of angles being randomly masked. The learning rate is adjusted based on the formula: $lr={lr}_{base}\times batch\,size\times GPU\,num\times grad\,accum\,steps$.
4.1.2 Dataset
Three datasets are used in our experiments, each containing 100k samples with a resolution of $512\times 512$:
Real-world dataset [8] consists of real sinogram images obtained from various materials and objects, derived from actual synchrotron radiation CT experiments, including Advanced Photon Source (APS [^2] Advanced Photon Source: https://www.aps.anl.gov) at ANL. This dataset, provided by the X-ray tomography data bank, includes a range of noise levels and artifacts that are typical in real-world settings, making it significantly more complex than simulated and synthetic datasets. Further, this dataset includes samples with dynamic features [28] and in situ [30] measurements, which captures wide-range of experiments at APS and other synchrotron facilities. This complexity offers a rigorous test of the model’s robustness and ability to handle practical challenges in sinogram inpainting.
Shape dataset generated using the scikit-image library [38] contains a variety of simulated geometric shapes, including circles, rectangles, and triangles. This dataset introduces a range of geometric complexities, allowing us to assess the model’s ability to reconstruct shapes with sharp edges and varying curvature.
Shepp2d dataset generated by the Tomopy library [13] comprises simulated shepp-logan phantoms, offering a controlled environment for evaluating the accuracy and stability of reconstruction algorithms. These phantoms are less complex and ideal for testing basic inpainting capabilities.
4.1.3 Metrics
We evaluate inpainting quality using Structural Similarity Index Measure (SSIM) [40] and Peak Signal-to-Noise Ratio (PSNR). SSIM quantifies structural preservation by measuring differences in structure, luminance, and contrast, while PSNR assesses reconstruction fidelity by comparing signal strength to noise levels. These metrics are chosen for their effectiveness in evaluating both structural accuracy and artifact reduction in sinogram restoration.
4.2 Accuracy Comparisons with Baselines
(a) Mask
(b) Ground Truth
(c) Ours
(d) UsiNet
(e) SinoTx
(f) StrDiffusion
(g) Copaint
Figure 3: Visual Results: Lines 1, 3, 5 show visual results of sinogram inpainting of the Real-world dataset, and lines 2, 4, 6 show the reconstructed images of the sinograms. Details of visual results for Shapes and Shepp2d datasets are in supplementary materials.
This section evaluates FCDM’s performance by comparing it against 9 baselines on 3 datasets. For each comparison, we focus on evaluating the SSIM and PSNR metrics specifically within the masked regions rather than across the entire image. This approach allows us to precisely measure the model’s ability to inpaint the most challenging portions of the sinogram, where data is missing or corrupted. Moreover, since the final goal of CT scans is reconstructed images, we also evaluate the images reconstructed by Gridrec[11] on the SSIM and PSNR metrics.
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.959 0.964 0.953 32.1 31.5 31.3
W/o Freq Conv 0.796 0.821 0.818 26.8 27.2 27.4
Spatial Conv 0.802 0.823 0.821 26.9 27.3 27.5
W/o $H$ Conv 0.855 0.859 0.857 26.5 26.8 27.7
W/o $W$ Conv 0.888 0.892 0.899 27.3 28.1 28.4
2D Freq Conv 0.899 0.906 0.910 27.8 28.3 28.7
Table 2: Ablation Results Related to Freq Conv: The experimental results show that the model’s performance is greatly impacted by the frequency domain convolution. Furthermore, there is a marginally more noticeable drop in performance when the $H$ dimension convolution is absent than when the $W$ dimension convolution is absent.
Tab. 1 summarizes the performance of our model and the baseline models on Real-world dataset [^3] Due to space limitation, please find our results on Shape and Shepp2d datasets in our supplementary materials.. In addition to the quantitative analysis, we also provide visual comparisons between the outputs of our model and the baseline models. Fig. 3 illustrates the inpainted sinograms and reconstructed images using Gridrec [11] when the mask ratio is equal to $0.8$, focusing on showing the visual quality and structural fidelity of the inpainted sinograms and reconstructed images. Our experimental results show that our model significantly outperforms all baselines on the more complex Real-world dataset. The simplicity of the Shapes and Shepp2d datasets allows previous baselines to achieve good results, making our model’s improvements relatively modest (but still obvious) in these cases. Real-world dataset presents greater challenges, and our model excels by effectively leveraging the unique properties of sinogram, leading to a clear and substantial advantage in inpainting quality over the other methods.
4.3 Ablation Studies
4.3.1 Impact of Bidirectional Frequency Domain Convolutions
We first examine the impact of BFDC by comparing several convolutional configurations. In addition to evaluating the performance of the full model with BFDC, we also test models that employ only spatial-domain convolutions, 2D frequency convolutions, and single-direction frequency convolutions along either the detector axis or the projection angles. Results show that removing frequency-domain convolutions leads to a performance drop, as purely spatial convolutions fail to disentangle overlapping sinogram features. While 2D frequency convolutions lack directional control and single-direction convolutions capture only partial spectral information, our bidirectional approach achieves the best balance between efficiency and feature preservation. We also compare peak memory allocation, runtime and FLOPs between FCDM incorporating frequency domain convolutions and without it, showing no significant difference (in Supplementary Materials).
4.3.2 Contribution of Tailored Loss Functions
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.959 0.964 0.953 32.1 31.5 31.3
W/o $L_{\text{adv}}$ 0.902 0.915 0.908 28.5 28.9 29.2
W/o $L_{absorp\_sum}$ 0.921 0.934 0.930 29.3 29.7 29.9
W/o $L_{freq}$ 0.927 0.938 0.929 29.8 29.6 30.1
Table 3: Ablation Results Related to Tailored Loss: The results illustrate the model’s performance on the real-world dataset when specific terms are removed from the loss function. It is evident that the model’s performance degrades when either $L_{absorp\_sum}$ or $L_{freq}$ is excluded.
The second part of our ablation study examines the role of the tailored loss functions, specifically the absorption consistency loss and the frequency domain loss. We conduct experiments where each of these losses is individually removed from the training process. Results in Tab. 3 demonstrate that the tailored losses are essential for achieving high-quality reconstructions. Similarly, the absence of the frequency domain loss results in less accurate inpainting.
4.3.3 Impact of Fourier-Enhanced Mask Embedding
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.959 0.964 0.953 32.1 31.5 31.3
W/o FEME 0.907 0.915 0.905 29.2 29.4 29.6
Table 4: Ablation results on FEME. Removing FEME reduces model performance, indicating its role in improving feature alignment.
To assess the contribution of FEME, we compare models trained with and without FEME in Tab. 4. The model trained with a standard 0/1 binary mask exhibits a drop in performance. By contrast, the FEME-enhanced model consistently achieves higher SSIM and PSNR scores, demonstrating that incorporating projection angle information helps the model better infer missing regions by leveraging angular dependencies.
4.3.4 Impact of Frequency-Adaptive Noise Scheduling
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.961 0.965 0.954 32.3 31.4 31.4
W/o Freq-Adaptive 0.912 0.923 0.916 28.9 29.1 29.4
Table 5: Ablation results on FANS. Without FANS leads to a decline in model performance, demonstrating its role in improving structural consistency.
Finally, we evaluate the impact of FANS by comparing two different noise injection strategies: standard Gaussian noise and our proposed FANS. As shown in Tab. 5, using standard Gaussian noise leads to lower SSIM and PSNR scores, as it does not account for the varying importance of different frequency components during the diffusion process. In contrast, our FANS improves performance, by ensuring early perturbation of high-frequency details while preserving low-frequency structures in later steps.
5 Conclusion
This work introduces FCDM, a new diffusion-based sinogram inpainting model that integrates bidirectional frequency-domain convolutions, Fourier-enhanced mask embedding, and frequency-adaptive noise scheduling. The evaluation results demonstrate that FCDM closely approximates the ground truth. Ablation studies further validate the effectiveness.
Supplementary Material
6 Complexity Analysis of Different Convolution Locations
6.1 Downsampling Followed by Frequency-Domain Convolution
Downsampling reduces the input size from $H\times W$ to $\frac{H}{r}\times\frac{W}{r}$, with complexity:
$\mathcal{O}_{down}=\mathcal{O}\left(\frac{2CWH}{r^{2}}\right).$ (23)
Half of the channels undergo standard convolution:
$\mathcal{O}_{s}=\mathcal{O}\left(C^{2}k_{s}^{2}\frac{WH}{r^{2}}\right),$ (24)
while the other half is processed via Fourier Transform, frequency convolution (Hadamard multiplication), and inverse Fourier Transform:
$\mathcal{O}_{f}=\mathcal{O}\left(C\frac{WH}{r^{2}}\log{\frac{WH}{r^{2}}}\right)+\mathcal{O}\left(C\frac{WH}{r^{2}}\right).$ (25)
A restore convolution follows:
$\mathcal{O}_{r}=\mathcal{O}\left(2C^{2}k_{r}^{2}\frac{WH}{r^{2}}\right).$ (26)
Finally, the result is processed by an Encoder $\mathcal{E}(\cdot)$ with complexity:
$\mathcal{O}_{e}=\mathcal{O}\left(\sum_{i=1}^{L}C_{i-1}C_{i}k_{i}^{2}\frac{H_{i}W_{i}}{r^{2}}\right).$ (27)
6.2 Latent Space Frequency-Domain Convolution
The input is encoded into a latent representation of size $C_{L}\times H_{L}\times W_{L}$:
$\mathcal{O}_{e}=\mathcal{O}\left(\sum_{i=1}^{L}C_{i-1}C_{i}k_{i}^{2}H_{i}W_{i}\right).$ (28)
Standard convolution is applied:
$\mathcal{O}_{s}=\mathcal{O}(C_{L}^{2}k_{s}^{2}H_{L}W_{L}).$ (29)
Frequency-domain convolution and inverse Fourier Transform follow:
$\mathcal{O}_{f}=\mathcal{O}(C_{L}H_{L}W_{L}\log{H_{L}W_{L}})+\mathcal{O}(C_{L}H_{L}W_{L}).$ (30)
A final channel restoration convolution is applied:
$\mathcal{O}_{r}=\mathcal{O}(2C_{L}Ck_{r}^{2}H_{L}W_{L}).$ (31)
7 Supplementary Evaluation Results
(a) Ours
(b) W/o Freq Conv
Figure 4: Visual Ablation Results: Results without frequency domain convolution are blurred.
Ours W/o Freq Conv
Memory (MB) 21505 21464
Time (s) 44.90 44.88
FLOPs (PMac) 1.07 1.07
Table 6: Computational cost comparison: Memory usage, execution time, and FLOPs differ by less than $1\%$, confirming that frequency-domain convolutions introduce negligible overhead.
Locations SSIM PSNR
Ds Followed by Freq Conv 0.962 33.9
Latent Freq Conv 0.987 35.9
Table 7: Ablation Results Related to Freq Conv Locations: Experiments are done in the first stage. Results show that by conducting frequency domain convolutions on the latent after the input has been passed through, an Encoder produces better performances.
(a) Mask
(b) Ground Truth
(c) Ours
(d) UsiNet
(e) SinoTx
(f) StrDiffusion
(g) CoPaint
Figure 5: Visual Results of Shape and Shepp2d Datasets: Line 1 and 3 are visual results of sinogram inpainting of Shape dataset and Shepp2d datset, respectively. Line 2 and 4 are visual results of the reconstructed images of the sinograms using iRadon[31]. The visual comparisons highlight our model’s ability to preserve structural details and reduce artifacts.
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.966 (0.924) 0.978 (0.938) 0.978 (0.955) 36.3 (34.8) 36.4 (35.3) 36.7 (34.3)
UsiNet 0.509 (0.480) 0.549 (0.511) 0.546 (0.512) 14.1 (10.9) 13.7 (12.6) 15.3 (12.8)
SionTx 0.746 (0.717) 0.743 (0.696) 0.796 (0.750) 26.5 (25.1) 26.9 (25.2) 27.1 (25.1)
StrDiffusion 0.524 (0.497) 0.552 (0.506) 0.600 (0.572) 14.3 (12.1) 15.1 (13.5) 15.9 (13.5)
CoPaint 0.903 (0.868) 0.929 (0.889) 0.908 (0.888) 30.2 (28.6) 31.0 (29.3) 33.0 (31.3)
CMT 0.454 (0.413) 0.469 (0.427) 0.477 (0.431) 19.9 (17.1) 20.2 (17.4) 21.0 (18.1)
hourglass 0.807 (0.772) 0.813 (0.781) 0.823 (0.781) 27.1 (24.7) 27.5 (24.3) 28.2 (26.5)
RePaint 0.915 (0.869) 0.918 (0.873) 0.926 (0.899) 33.1 (32.3) 33.6 (31.7) 33.69 (31.2)
MISF 0.823 (0.798) 0.847 (0.825) 0.862 (0.842) 30.2 (28.9) 30.7 (29.4) 31.1 (29.8)
LaMa 0.852 (0.821) 0.868 (0.835) 0.885 (0.831) 30.5 (29.1) 31.2 (30.0) 31.7 (30.3)
CoIL 0.746 (0.683) 0.786 (0.709) 0.809 (0.734) 27.9 (25.3) 28.1 (25.7) 29.0 (26.6)
Table 8: Quantitative Accuracy Results on the Shape Dataset: The results demonstrate that our model excels in handling sinograms of objects with sharp edges and geometric structures, outperforming the baselines in the inpainting performance.
Methods SSIM PSNR
Mask Ratio Mask Ratio
0.4 0.6 0.8 0.4 0.6 0.8
Ours 0.967 (0.937) 0.979 (0.955) 0.978 (0.962) 37.8 (36.7) 37.3 (34.6) 37.9 (35.6)
UsiNet 0.494 (0.447) 0.499 (0.462) 0.492 (0.452) 13.1 (12.1) 15.0 (13.2) 15.6 (14.5)
SionTx 0.967 (0.938) 0.968 (0.926) 0.962 (0.906) 37.2 (35.8) 37.5 (34.4) 36.3 (33.9)
StrDiffusion 0.516 (0.497) 0.554 (0.505) 0.562 (0.538) 15.5 (12.5) 16.3 (14.9) 16.7 (15.3)
CoPaint 0.932 (0.892) 0.947 (0.891) 0.946 (0.914) 32.0 (30.7) 31.8 (30.5) 32.7 (32.0)
CMT 0.504 (0.457) 0.507 (0.477) 0.518 (0.485) 20.8 (19.3) 22.6 (19.3) 22.7 (19.5)
hourglass 0.826 (0.770) 0.843 (0.813) 0.851 (0.818) 27.5 (24.9) 27.7 (26.3) 28.1 (26.9)
RePaint 0.912 (0.877) 0.931 (0.894) 0.934 (0.913) 33.2 (31.3) 34.3 (31.2) 35.2 (33.5)
MISF 0.832 (0.812) 0.849 (0.826) 0.843 (0.821) 31.5 (30.2) 32.1 (30.8) 32.8 (31.4)
LaMa 0.821 (0.798) 0.834 (0.812) 0.847 (0.823) 31.2 (30.0) 31.9 (30.4) 33.1 (31.7)
CoIL 0.814 (0.749) 0.855 (0.795) 0.870 (0.809) 28.9 (27.8) 29.1 (28.0) 30.1 (28.8)
Table 9: Quantitative Accuracy Results on the Shape2d Dataset: The controlled and simplified nature of the Shepp2d dataset enables previous baselines to perform well, yet our model still demonstrates clear improvements in the inpainting performance.
7.1 Supplementary Comparison Results
Tab. 8Tab. 9, and Fig. 5 present the quantitative and qualitative comparisons on the Shape and Shepp2d datasets against several state-of-the-art baselines. For both datasets, FCDM consistently achieves higher SSIM and PSNR scores compared to previous baselines, indicating superior inpainting accuracy of our model.
On the Shape dataset, which consists of geometric shapes with distinct edges and simple textures, our model demonstrates strong performance in accurately completing sinograms of objects with sharp boundary features. For the Shepp2d dataset, despite its simpler structure, the results confirm that our model can still effectively capture its features and maintain high-quality inpainting. Although the improvements in these controlled datasets are relatively modest, they showcase our model’s robustness and consistency in diverse scenarios, demonstrating that it can perform on par with or better than previous baselines.
Fig. 4 provides visual results from the ablation studies conducted on the Real-world dataset, offering a comprehensive analysis of the contributions of frequency-domain convolutions. The visual comparisons clearly illustrate that removing the frequency-domain convolutions leads to notable degradation in inpainting quality.
7.2 Memory Usage Caused by Frequency Domain Convolution
We conduct an additional experiment to assess the impact of frequency-domain convolutions on computational efficiency by measuring peak memory allocation, inference time, and FLOPs. Results in Tab. 6 indicate that incorporating frequency-domain convolutions introduces less than a 1% increase in memory usage and execution time, with no significant change in FLOPs.
This demonstrates that performing frequency-domain convolutions in the latent space minimizes overhead while preserving inpainting accuracy. Unlike applying convolutions on full-resolution images, this approach reduces computational complexity while effectively capturing structural information, making it feasible for large-scale sinogram inpainting.
7.3 Effect of Frequency-Domain Convolution Placement
We compare the placement of frequency-domain convolutions before and after the Encoder within the Encoder-Decoder component. Results in Tab. 7 show that applying frequency convolutions after encoding, in the latent space, achieves better inpainting quality.
This improvement is likely due to the better abstraction and compression achieved by the Encoder, which filters out redundant noise before frequency processing. In contrast, applying frequency convolutions before encoding may introduce spectral distortions, reducing reconstruction accuracy. Processing in the latent space enables more effective feature extraction, improving structural consistency across projection angles.